### v1

Abstract
How informative are job ads about the actual pay and amenities offered by employers? Using a comprehensive database of job ads posted by Norwegian employers, we develop a methodology to systematically classify the information on both pay and non-pay job attributes advertised in vacancy texts. We link this information to measures of employer attractiveness, which we derive from a job search model estimated on observed wages and worker mobility flows. About 55 percent of job ads provide information related to pay and nearly all ads feature information on non-pay attributes. We show that publicly advertised job attributes are meaningful predictors of employer attractiveness, and non-pay attributes are about as predictive as pay-related attributes. High-pay employers mention pay-related attributes more often, while high-amenity employers are more likely to advertise flexible working hours and contract duration.
1 Introduction
There is a long tradition in labor economics of seeing jobs as more than pay (e.g., Brown, 1980; Duncan and Holmlund, 1983; Rosen, 1986). Expressions commonly used in the literature such as “compensating differentials,” “amenities,” or “workplace differentiation” all conceptualize the notion that jobs can be represented as a multidimensional bundle of attributes that workers trade off when considering opportunities with alternative employers. Such trade-offs can be substantial. Recent experimental evidence shows that workers at a US call center are willing to give up 8 percent of their salary for the option to work from home (Mas and Pallais, 2017), while undergraduates at a top US university would accept a 5.1 percent salary cut for a job offering the possibility to work part-time (Wiswall and Zafar, 2018).
What is unclear from this evidence, however, is how workers gather information on job attributes—both pay and non-pay—when exploring their options in the labor market. One important search channel through which workers can learn about these attributes is the content of job postings. [^4] Recently, Carrillo-Tudela et al. (2023) provide survey evidence from Germany showing that 55.3 percent of firms used job postings in their latest hire, while 88.1 percent of workers used job postings in their search process. Notably, however, job postings are not exclusive of other search methods, such as social networks and recruitment agencies. Carrillo-Tudela et al. (2023) find that firms and workers use, respectively, 1.9 and 2.3 channels on average. As employers ultimately decide on the content that they wish to mention in publicly advertised job postings, it remains an open question whether such content in fact reflects the actual pay and amenities associated with different employers.
In this paper, we study in detail the information content of job ads. Our first objective is to systematically quantify what workers can learn about the pay and non-pay attributes of jobs from the texts of job ads. Our second objective is to provide a detailed description of how different employers–characterized both in terms of their pay premiums and amenity values–advertise different types of pay and non-pay attributes in job ads. Our third objective is to provide statistical measures of how well publicly advertised job attributes predict employer pay premiums and amenity values. Our fourth and final objective is to introduce quantitative measures of ad informativeness in an economic framework that allows us to study counterfactual job mobility flows under alternative informational environments.
To characterize the distribution of job attributes advertised by employers in job ads, we use the near-universe of publicly posted vacancies in Norway between 2002 and 2019. A key contribution of our paper is to systematically extract job attributes from these texts. Using tools from natural language processing, we identify close to fifty commonly advertised pay and non-pay attributes, such as “competitive pay,” “flexible work hours,” and “nice work environment.” We build a large collection of unique expressions associated with each attribute, which we make publicly available online for other researchers to use. [^5] The full list of expressions is available in an online repository, accessible through this link. This approach allows us to identify a large number of job attributes and to transparently map each attribute to many alternative expressions.
We find that about 55 percent of job ads provide some pay-related information, with 27 percent revealing explicit information on salaries, such as by mentioning an actual salary number, a salary bracket, or indicating that pay is set by a collective agreement. Moreover, nearly all ads feature information on some amenity attributes, such as contract duration, irregular hours, shift work, flexible work hours, workplace attributes, task-related attributes, or other minor perks. These results clearly suggest that employers use the texts of publicly posted job ads to advertise characteristics of jobs that workers potentially value.
A key strength of our data set is that each ad can be directly traced to the establishment posting the job. To shed light on how different employers advertise different types of information in publicly posted job ads and to assess the information content of these publicly advertised job attributes, we estimate workers’ valuation of employers using job flows and wages observed in matched employer-employee data. Building on prior work by Sorkin (2018) and Morchio and Moser (2024), we use a structural revealed preference approach and obtain measures of workers’ valuation of different employers. Within our framework, this valuation can be further decomposed into a pay component, a non-pay component, and a job security component. Measuring these components separately matters because many of the attributes advertised in job listings, such as workplace quality, are not necessarily reflected in pay.
By linking information on publicly advertised attributes in job ads to the employer-level estimates from our structural framework, we can measure the type of attributes posted by different employers. We find that high-pay employers are more likely to advertise compensation-related information, while high-amenity employers more often mention job amenities, such as “possibility to work flexible hours” and “regular daytime work schedule”. This systematic relationship between workers’ estimated valuation of employers, which is derived from actual labor market outcomes, and the propensity of these employers to mention specific attributes is evidence that employers supply credible pay and non-pay content in their job listings.
We show that the content of job ads has predictive power for workers’ valuation of employers. Notably, we find that the attributes we detect in job ads can explain about 60 percent of the variation in overall employer values and employer pay premiums, and about 50 percent of non-pay and 40 percent of job security differences across employers. Interestingly, job ads have predictive power over and above commonly observed job characteristics. Once we remove variation in employer values that can be attributed to flexible industry, occupation, and location controls, job attributes listed in ads can still explain up to 20 percent of the remaining variation across employers. In this regard, we find that attributes related to having flexible working hours and shift work are especially important predictors. Further, controlling for industry, occupation and location, we show that employers with higher posted pay premiums and amenity values in job ads have shorter average duration of posted vacancies, consistent with the notion that such employers are more attractive.
Finally, we use our model estimates to gauge the impact on worker mobility of having access to the content of job ads. We predict the counterfactual value of employers under alternative assumptions on workers’ information sets and derive the implied mobility flows in each case. In our baseline, we find that aggregate job-to-job transitions to more highly-valued employers increase by 1.6 percent in an economy where workers have access to information on job attributes in job ads, as compared to a benchmark economy where only “one-liner” vacancies listing occupation, industry, and location are posted. This increase corresponds to 25 percent of the additional mobility towards better employers in the full-information model. Our results further suggest that the information in job ads can contribute to men and women self-selecting into distinct employers based on their preferences.
We see our approach as the first attempt at quantifying the breadth and quality of information present in the texts of job ads. We show that job openings contain a wide range of attributes relevant to workers’ valuation of employers and that a subset of these attributes aligns with these valuations: the content of job ads is informative about the actual attractiveness of employers. These findings have implications for a wide range of models in the job search literature, such as models with a multi-dimensional job ladder (Hwang et al., 1998; Jarosch, 2023). We leave it to future research to shed light on employers’ endogenous choices of providing more or less content in their job ads, and the type of content they provide, in relation to their recruitment objectives.
Our results also have implications for the design of policies related to the information contained in job ads. There is a current policy drive to improve pay transparency within organisations in many countries, and several jurisdictions are also implementing policies to promote pay transparency in job openings Cullen (2024). [^6] As an example, in November 2022 New York City mandated employers to systematically include a salary bracket in their job ads. Taken together, our results on the prevalence of specific job attributes in vacancy texts and on the type of attributes advertised by better employers can be helpful to inform further policies in that direction.
Related literature
Several recent studies have used the job description of vacancies as a source of information on the type of jobs advertised by employers. A large portion of this literature centers on skill requirements. Marinescu and Wolthoff (2020) focus on the role of job titles in accounting for the number of applicants received by each vacancy. Deming and Kahn (2018), Atalay et al. (2020), and Deming and Noray (2020) study variations in the demand for specific skills across local labor markets and over time using information extracted from the job description of vacancies. Many recent contributions specifically focus on the demand for computer science and artificial intelligence skills (see, e.g., Alekseeva et al., 2021; Acemoglu et al., 2022; Braxton and Taska, 2023; Contractor and Taska, 2022).
We contribute to this literature by instead considering a broad set of pay and non-pay attributes that are advertised by employers in job listings. In a similar vein, several recent contributions focus on a single specific attribute, such as the type of flexible work arrangements advertised in job ads (Adams-Prassl et al., 2020, 2023a), the possibility to work remotely (Hansen et al., 2023), and whether the employer offers training on the job (Adams-Prassl et al., 2023b). We add to these important contributions by retrieving a comprehensive set of job attributes advertised in the text of job openings, which encompasses flexible work arrangements, remote work and on-the-job training, but also what employers communicate about compensation or the quality of the workplace in their vacancies. In a similar spirit, Sockin (2022) uses text analysis to retrieve information on a large set of amenities from employer reviews posted by workers, rather than using vacancy texts. We see our results using vacancy texts as complementary since they instead reflect the information advertised by employers. [^7] In other recent and related work, Arold et al. (2024) and Lagos (2024) use text analysis methods to retrieve information on amenities contained in collective bargaining agreement texts.
A large body of theoretical and empirical work has emphasized that non-pay attributes are important determinants of how much workers value alternative employers. One strand of this literature elicits workers’ preferences over specific attributes in survey experiments or quasi-experimental settings. Examples of non-pay attributes analyzed using this approach include workplace flexibility (Mas and Pallais, 2017; Wiswall and Zafar, 2018; Drake et al., 2023), job security (Datta, 2019), commuting time (Le Barbanchon et al., 2021), working from home (Nagler et al., 2022; Lewandowski et al., 2022), and shift work (Desiere and Walter, 2023), among others. Our paper instead provides a composite measure of employer amenity value that is linked to a large number of non-pay job attributes from job ads.
We follow the second strand of this literature, which recovers structural estimates of the non-pay value of employers by modelling where workers choose to work in a labor market with heterogeneous employers. In this approach, estimates of the non-pay value of employers are obtained by reconciling workers’ choice of employer with their level of pay (Rosen, 1986). The choice of workers are captured using either the size of employers in a static setting (e.g., Card et al., 2018; Lamadon et al., 2022) or mobility flows between employers (e.g., Sorkin, 2018; Taber and Vejlin, 2020; Morchio and Moser, 2024). We rely on the latter approach since it can be recast in a job search framework, allowing us to further disentangle the amenity value of employers from differences in their recruitment effort. Relative to these important studies, we link the estimates of workers’ valuation of employers to the attributes advertised in job openings and quantify the value of this information. [^8] Morchio and Moser (2024) also relate their estimates of the amenity values of employers to some of the workplace characteristics observed in their data, such as “workplace hazards” and “working hours flexibility”.
We also contribute to the expanding body of work that aims at unpacking search frictions in the labor market. Jäger et al. (forthcoming) elicit the beliefs of workers on their reemployment wage if they were to lose their job and compare these wages to actual labor market outcomes. Horton et al. (2024) and Belot et al. (2022) design experiments within a controlled environment to study how job seekers respond to variation in the wage featured in job vacancies. Our approach adds to these studies in that we focus on the texts of actual job postings as a source of information on alternative employers. Our evidence also relates to recent studies on how wages relate to the duration of posted vacancies (e.g., Mueller et al., 2024; Bassier et al., 2023; Faberman and Menzio, 2018). [^9] Several recent studies have examined the wage elasticity of job applications (e.g., Azar et al., 2022; Banfi and Villena-Roldan, 2019) and recruitment (e.g., Bassier et al., 2022; Hirsch et al., 2022). We add to this literature by showing how measures of posted pay premiums and amenities relate to vacancy duration.
Outline
The rest of the paper is organized as follows. Section 2 describes the data. Section 3 explains how we retrieve information on the attributes advertised by employers in vacancy texts and provide evidence on their prevalence in our data. Section 4 introduces our structural model and describes the estimated values of employers. Section 5 provides our evidence based on the linking of information in job ads to our structural estimates and reports our results on the information content of the texts of job ads. Section 6 concludes.
2 Data and Institutional Context
2.1 Vacancy Data
We have access to almost 3 million job ads covering the near universe of publicly posted vacancies in Norway between 2002 and 2019. These data are maintained by the Norwegian Public Employment Agency (NAV). The employment agency collects information about vacancies from several sources, including online job boards and newspapers, as well as vacancies that are directly reported by employers to the agency.101010 In accordance with the Labor Market Act §7, Norwegian employers are required by law to report publicly posted job ads to the agency, which maintains a comprehensive database of publicly posted vacancy ads with the stated goal of providing job-seekers with current information on suitable job opportunities. More than 80 percent of job postings recorded in the agency’s database were retrieved from various online job boards. [^11] Bhuller et al. (2023) study the consequences on labor market matching of increased online job search and recruitment triggered by a roll-out of broadband internet across Norway during the early 2000s. The remaining 20 percent were either (i) scanned or transcribed from newspapers by caseworkers in the employment agency, or (ii) based on notifications of job openings sent by employers directly to the agency. The ability to observe virtually all publicly posted vacancy ads in the economy with full-text corpora is an important advantage of our setting, which differentiates it from the existing literature using vacancy texts that often relies on information from selected online job portals, typically covering job ads posted over a shorter time span.
In addition to the job title and actual text of each job opening, our data contain the following structured information about each ad: unique establishment identifiers, the dates when the ad was registered and filled or removed (i.e., vacancy duration), the number of job openings per vacancy, and some additional information about job characteristics submitted by employers. The establishment identifiers are central to our analysis. They allow us to link each job opening to matched employer-employee administrative data and to compare the information we extract from the text to actual outcomes at the establishment level.121212 Notably, around 7 percent of job ads in our data were posted through recruitment or temporary employment agencies, and we drop these from most of our analysis, as we are unable to link such ads to the actual establishment where the job is placed. Further, in around 13 percent of job ads, the posting employer has for various reasons decided not to disclose the establishment name in the publicly posted information, but we do have the corresponding establishment identifier in our data as the agency maintains this information and could share this with us for research purposes (by the Labor Market Act §7-4). Each job ad in our data set also has an occupational code based on the 4-digit ISCO classification. This occupational code is assigned by caseworkers based on the job title and textual information on job descriptions and skill requirements stated by the employers in the vacancy, and is a novel feature of our data. Besides the structured information, the text of a job ad typically contains about 200 words. As we describe in Section 3, we use tools from the natural language processing literature to retrieve information on job attributes from these texts.
2.2 Matched Employer-Employee Data
We also have access to Norwegian administrative matched employer-employee data for all years between 2000 and 2019. The data come as a series of employment spells in each year, where each spell has information on the individual’s employer, as well as their pre-tax earnings, hours of work, and employment start and end date. Individuals can have several recorded spells in any given month if they receive earnings from more than one employer. For each worker, we aggregate all employment spells to the annual level and retain observations corresponding to the main employer, defined as the establishment with the largest annual earnings. Using information on annual earnings and annual contracted hours of work, we further calculate the average hourly wage for each worker in their main job in each year.
This data set has additional information on several background characteristics of employees (gender, education) and employers (location, industry). In the most recent years, the data set also has some information on a small subset of job attributes (e.g., shift work, contract type) for each employment spell that employers must report to Statistics Norway.
2.3 Sample Selection and Analysis Periods
While our vacancy and employer-employee data sets cover almost two decades, both in our descriptive and structural analyses we split the data into four five-year periods: 2000-2004, 2005-2009, 2010-2014, and 2015-2019. We do this in order to capture potential changes in workers’ valuation of employers in terms of their pay and non-pay attributes, as well as employers’ posting behavior. In most of our analysis, we focus on the 2015-2019 period as our baseline and provide supplementary results for the remaining five-year periods.
We further impose several restrictions that limit our focus to establishments for which we observe both job ads and job-to-job flows. Table 1 describes these restrictions and shows how they reduce the number of observations used in our baseline analysis for the 2015-2019 period. [^13] Appendix Table A.1 documents these sample restrictions for each of the four periods used in the analysis, as well as the 2000-2019 period as a whole. Our initial matched employer-employee data contains about 300,000 establishments covering about 3.1 million unique workers and 13.3 million worker-year observations. During the 2015-2019 period, we observe about 900,000 job ads. [^14] An ad is on average associated with 1.6 job openings, so these ads cover 1.4 million job openings. First, we limit our analysis to job ads with complete vacancy texts and to establishments that posted at least one job ad during the corresponding five-year period. [^15] As we explain in Section 4, we also group similar establishments using a k-means clustering approach. Appendix Table A.2 shows that restricting the sample strictly to establishments with observed job ad texts would only slightly lower our sample sizes. However, using the clustering approach greatly reduces the dimension of the number of employer-specific parameters we need to estimate in our structural analysis. For almost 60 percent of establishments observed in our employer-employee data, we do not observe any job ad texts in the corresponding five-year period. Despite this restriction, the remaining establishments cover almost 90 percent of worker-year observations, reflecting that posting establishments are larger. [^16] See Appendix Table A.3 for a comparison of establishment characteristics for different sample restrictions. Second, we remove job ads posted by staffing agencies, as for such ads we observe the establishment identifier for the staffing agency and not the actual employer. [^17] Around 92.5 percent of job ads classified as “staffing agency” ads in our data were posted by temporary employment agencies (e.g., Manpower), while recruitment agencies account for the remaining share. Third, our structural estimation is further restricted to establishments in the set of strongly connected establishments, i.e., the set of establishments across which we observe two-way job mobility flows during the corresponding five-year period. This restriction alone excludes almost 75 percent of establishments observed in our employer-employee data, but more than 80 percent of worker-year observations are still retained. Lastly, when we relate the employer-level model estimates to the information extracted from job ads, we impose both restrictions simultaneously. Interestingly, in this sample, we still observe job ads for nearly all establishments in the strongly connected set.
Table 1: Overview of Sample Selection: Baseline Analysis Period.
Job Ads Establishments Workers Worker-Years
(1) (2) (3) (4)
All Observations $915,790$ (100%) $301,716$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $127,913$ (42.4%) $2,733,050$ (87.5%) $11,898,501$ (89.6%)
Excluding Staffing Agencies $774,076$ (84.5%) $127,904$ (42.4%) $2,732,297$ (87.5%) $11,895,973$ (89.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $78,118$ (25.9%) $2,499,506$ (80.0%) $10,974,896$ (82.7%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for the 2015-2019 baseline period. “Observe Vacancy Text” is the set of vacancies with non-missing text content and at least one section written in Norwegian. “Excluding Staffing Agencies” is the subset of ads with observed text that are not posted by recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural estimation. Sample sizes are calculated after linking job ads to matched employer-employee data at the establishment-cluster level. Appendix Tables A.1-A.2 provide sample selection for each of the four five-year periods (2000-2004, 2005-2009, 2010-2014, and 2015-2019) and the combined 2000-2019 period, with linking job ads to matched employer-employee data at the cluster and the establishment level, respectively.
2.4 Job Postings and Recruitment
An underlying assumption in our analysis is that job postings are an important source through which employers recruit workers and where workers learn about available jobs on the market. In order to assess the importance of job postings for recruitment, we link each ad in our vacancy data to a potential hire observed in our employer-employee data. Building this linkage requires us to assign a potential hire to each job ad by finding a job spell that starts (i) in the same establishment as the job posting, (ii) soon after the ad is posted, and (iii) in the same or a close occupation as the one assigned to the vacancy. We are able to link 89.4 percent of all publicly posted non-staffing agency job ads to a hire within the following six months; this applies to 94 percent for ads posted by public sector state employers, 90.2 percent of ads by non-state public sector employers (e.g., local municipalities, public schools, public hospitals), and 87.2 percent of ads by private sector employers. Conversely, 56.6 percent of all hires can be related to a posted ad within the previous six months; this applies to 90.1 percent of hires by public sector state employers, 62.8 percent of hires by non-state public employers, and 48 percent of private sector hires. [^18] While all employers have an obligation to report the vacancies they post to the Norwegian employment agency, it is up to them to decide whether or not to post a vacancy for the purposes of recruitment. For instance, employers may also use other recruitment channels, such as social networks, or directly solicit applications. However, recruitment by public sector state employers is regulated by state legislation, which requires as a general rule that all job openings are publicly advertised (the State Employees Act §4). This evidence suggests that job postings are indeed an important channel for job search and recruitment in the Norwegian labor market.
3 Pay and Non-Pay Attributes in Vacancy Texts
This section describes how we retrieve the job attributes publicly advertised by employers in job ads. These attributes are extracted from vacancy texts using tools developed in natural language processing. To the best of our knowledge, we are among the first to systematically extract a comprehensive set of job attributes from the texts of job ads.
3.1 Extracting Advertised Job Attributes
Consider the sample vacancy text for an IT-consultant position in Figure 1. This text contains a variety of information on the type of tasks associated with the position, the required skills and experience necessary to perform these tasks, and the pay and non-pay attributes of the job. As this example makes clear, these attributes cover many different aspects of the position, ranging from the duration of the contract and regular hours (“full time, permanent position”) to characteristics of the workplace (“good working environment”) and information about the level of pay (“competitive conditions”). Our goal is to systematically extract this information from the open text of all vacancies in the database.
Figure 1: Sample Vacancy Text: IT-consultant.
IT-consultant About the job: Full-time, Permanent position We are looking for a new team player for our IT department! As an IT consultant, you will be responsible for our IT environment. It also includes getting to know large parts of our organization well through having internal user support as well as employee training. You simply get a varied working day, with different tasks to brush up on. In the future, you will also assist in larger IT development projects […] Tasks: - IT security. - SharePoint. - Data warehouse. - SOX Compliance. Required qualifications: - Likes IT technical challenges. - Sociable and likes to communicate with people. - Has relevant education. - Has experience from similar positions. We can offer: - A workplace with many varied tasks and a high focus on quality. - Training in our key products/services. - Good working environment with solid professional expertise. - Flexible working hours. - Competitive conditions. […]
Note: Translation from Norwegian by the authors. In Appendix Figures A.1-A.2, we provide additional examples of representative vacancy texts covering a teaching substitute and a civil engineer, respectively.
The key difficulty with extracting job attributes from the raw texts of job ads is that a single attribute can be expressed in many distinct ways in natural language. Besides, while prior work on the pay and non-pay attributes of jobs has centered on specific amenities, we do not have a definitive list of the job attributes advertised by employers in vacancy texts. We therefore proceed in three broad steps to extract these attributes from all job postings: (i) we pin down a list of attributes to extract, (ii) we ascribe a set of expressions to each of these attributes, and (iii) we label the entire corpus of vacancy texts based on these expressions.
Step 1: List of job attributes to include
Our choice of which pay and non-pay job attributes to include in the analysis is based partly on information that is commonly stated in ads, partly on our knowledge of the particular context of the Norwegian labor market, and partly on attributes that are considered important in the existing literature.
We identify commonly mentioned attributes in publicly posted vacancies using two distinct strategies: human recognition on a subset of job ads in our data set and identifying commonly used phrases in the “we offer” sections of vacancy texts. We first had several research assistants read through the text of 1,200 randomly chosen job ads in the vacancy corpus. The research assistants were asked to make a list of the commonly advertised attributes that are valuable to workers, to take note of the corresponding expressions associated with these attributes, and to indicate the presence of these attributes in each job posting.
As a second source of information on commonly advertised attributes, we isolate and extract common phrases from the “we offer” sections of job ads. We do this in two steps: First, we identify and extract all structured lists from the vacancy corpus. [^19] These lists are found by searching for consecutive sentences starting with a hyphen or bullet point and by searching for the HTML tags used to generate lists in online vacancies. By design, this approach identifies attributes that can be organised in structured lists. As an example, in the sample vacancy text in Figure 1, we retrieve three distinct lists: “Tasks,” “Required qualifications,” and “We can offer.” We find that 50 percent of the vacancy texts in our sample feature one or more such lists. Second, we isolate “we offer” lists by applying unsupervised topic modeling (Blei et al., 2003) to the collection of lists retrieved in the previous step. Further methodological details are provided in Appendix B. The most common expressions in the “we offer” lists constitute the second source of information on job attributes that we are able to identify.
Based on the two approaches described above, we are able to identify 47 distinct job attributes, as well as an initial set of expressions associated with each attribute. We summarize these 47 attributes in ten broad categories within classes of pay and non-pay job attributes, as shown in Table 2. Taken together, these attributes cover a variety of different job characteristics advertised in the text of job ads, such as information on career opportunities, convenient/inconvenient hours, and the quality of the workplace. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
Table 2: Main Categories of Job Attributes.
Main Category Examples
A. Pay Attributes
   – Compensation Scheme Salary band; Bonuses; CBA-based pay
   – Financial Attributes Insurance scheme; Pension scheme
   – Career Opportunities On-the-job training; Career opportunities
B. Non-Pay Attributes
   – Hours of Work Full-time; Part-time; On-call job
   – Convenient Hours Regular daytime work; Flexible working hours
   – Inconvenient Hours Shift-work; Weekend/evening/nights
   – Contract Duration Permanent job; Temporary job
   – Workplace Attributes Social environment; Remote work
   – Task-Related Attributes Interesting tasks; Challenging tasks
   – Other Minor Perks Location; Company vehicle; Company gym
Note: This table documents broad categories of attributes detected in vacancies, and examples of individual attributes. The full list of job attributes with descriptions is provided in Appendix Tables B.3-B.5.
Step 2: Expressions corresponding to each job attribute
Each job attribute retrieved in Step 1 is associated with an initial set of expressions. For instance, the attribute “flexible working hours,” is associated with the target expressions “flexible working hours” and “flexible work time arrangements.” [^20] The corresponding expressions in Norwegian are “fleksibel arbeidstid” and “fleksible arbeidstidsordninger,” which denote working hour schemes that offer workers the possibility to choose their own schedule. In Step 2, we seek to enrich this initial set by including other common ways of referring to the same attributes.
To expand on the set of possible expressions in our analysis, we rely on the Continuous Bag-of-Words (CBOW) Word2Vec algorithm (Mikolov et al., 2013). [^21] See Atalay et al. (2020) for an application of this algorithm to vacancy job titles. This approach measures the degree of similarity between phrases by exploiting the fact that words with similar meanings tend to occur in similar contexts. As an example, information about flexible working hours is commonly given with information about whether the spell is full-time or part-time in our corpus. We train the CBOW model using our collection of job ads, thus associating words that are similar in the context of job ads but not necessarily in other contexts. We then loop through the full collection of job ads and use the trained model to store phrases of one, two, or three consecutive words most similar to the phrases in our original lists from Step 1. Finally, we remove expressions that are wrongly classified as similar by the model. An example of a removed phrase is “flexible job adaptation,” which is classified as similar to “flexible working hours” by the model, but does not imply flexibility in one’s schedule. [^22] Notably, as there are two variants of written Norwegian (Bokmål and Nynorsk), we translate all phrases from Bokmål to Nynorsk, and append phrases that are not already captured by our dictionary. This discussion highlights the main advantage of our approach. We are able to identify a large number of attributes that can be described by many different expressions, but we still retain control on which specific expressions are included.
In total, our dictionary consists of 1,772 unique expressions, and each of the 47 attributes is associated with approximately 39 expressions on average, although some attributes have many fewer associated expressions than others. We have made the full list of expressions is available in an online repository, which is accessible through this link. Differences in the number of associated phrases reflect both the number of ways an attribute is presented in vacancy texts and the specificity of an attribute. For instance, the attribute “inclusive work-life scheme” is usually discussed using specific terms in Norwegian, which is why the model identifies only a small number of expressions for this attribute. Conversely, the attribute “shift work” is associated with many alternative expressions since there are many ways jobs can involve shift work.
Step 3: Apply to all vacancies
We simply use our 1,772 unique expressions to generate job attribute indicators for all ads in the data. A job ad is defined as advertising a given attribute if any of the expressions associated with that attribute is found in the job description.
3.2 Validation
A potential drawback of our approach is that it might fail to detect attributes implied by more complex phrases. Consider the sentence “we offer flexibility in starting date as well as in working hours.” As there is some text between “flexibility” and “working hours,” our dictionary approach would not detect flexibility in working hours in this case. More generally, we can expect that a job posting does offer a specific attribute when our procedure identifies it (high precision rate), but we can be less certain that a job posting does not contain that same attribute when it is not identified by our procedure (low sensitivity rate).
Table 3: Validation of Detected Job Attributes.
All Ads Job Ads in the Validation Sample
Text Analysis Text Analysis Manual Recognition Success Rate Precision Rate Sensitivity Rate
(1) (2) (3) (4) (5) (6)
A. Pay Attributes 67.1 79.2 73.2 89.0 89.3 96.6
   –– Compensation Scheme 53.1 64.8 61.8 92.5 91.9 96.4
   –– Financial Attributes 30.4 37.8 34.2 96.5 90.7 100.0
   –– Career Opportunities 41.1 48.8 41.8 79.0 71.3 83.2
B. Non-Pay Attributes 95.8 98.0 97.0 97.5 98.2 99.2
   –– Hours of Work 76.9 77.2 72.2 92.0 91.6 97.9
   –– Convenient Hours 20.3 15.8 15.8 94.0 81.0 81.0
   –– Inconvenient Hours 24.3 26.5 21.5 93.5 78.3 96.5
   –– Contract Duration 63.6 67.2 66.8 88.0 90.7 91.4
   –– Workplace Attributes 47.7 55.0 51.0 82.0 80.0 86.3
   –– Task-Related Attributes 52.6 65.0 46.5 72.0 64.2 89.8
   –– Minor Perks 39.5 46.0 18.5 66.5 33.7 83.8
Any Observed Attribute 96.4 98.2 98.0 98.2 99.0 99.2
Average Number of Attributes 5.84 6.64 5.56
Number of Job Ads 915,790 400 400 400
Note: This table compares prevalence rates for ten broad categories of job attributes measured using our text analysis approach and manual recognition. A broad category of job attribute is considered present if at least one of the underlying distinct job attributes is detected. Column (1) is for the full sample of job ads posted between 2015 and 2019, while Columns (2)-(6) consider a random sample of 400 job ads used in the validation. Columns (1)-(3) report prevalence rates, while Columns (4)-(6) provide summary statistics that compare prevalence rates from text analysis and manual recognition. Success rate is defined as the share of job ads where the text analysis yields the same result as manual recognition. Precision is the rate of agreement between the two methods given that an attribute was detected in our text analysis, while, sensitivity is the rate of agreement given that an attribute was detected in the manual recognition. Precision is a measure of how many detected attributes are false positives, and sensitivity is a measure of how good our method is at recovering true attributes. See Appendix Table A.4 for results from validation exercises for each of the 47 underlying job attributes that contribute to the ten broad categories shown here.
To check how important this limitation is in our setting, we performed an additional round of manual recognition. We selected another random sample of 400 job ads posted during the last five-year period, i.e., 2015-2019, and asked another group of research assistants to manually classify job attributes in this sample. We deliberately recruited a distinct group of assistants from the ones who helped us at the initial manual classification step (Step 1 of Section 3.1). We gave each research assistant our full list of 47 attributes and a few sample expressions for each attribute, and they were tasked with searching for these attributes in the full text of each job ad. We can then systematically compare the attributes retrieved with our automated procedure to those detected in the manual recognition performed by the assistants.
Table 3 shows the results of this comparison for ten broad categories of job attributes. Column (1) reports the prevalence rates measured based on our text analysis for each of these categories in the full sample of job ads from 2015–2019 used in our analysis. Column (2) reports the corresponding rates for the random sample of 400 job ads, also based on our text analysis procedure, while Column (3) reports the prevalence rates from manual recognition for the random sample. Finally, Columns (4)-(6) provide rates of success, precision, and sensitivity by comparing prevalence rates for each category across text analysis and manual recognition.
Across the ten broad categories of job attributes, we find that our procedure performs well in terms of success, precision, and sensitivity. Success and precision rates are above or around 80 percent for most categories. The notable exceptions are “task-related attributes”, where we find a success rate of 72 percent and precision of 64.2 percent, and “minor perks”, where we find a success rate of 66.5 percent and precision of 33.7 percent. The former is driven by differences in detection rates for whether the job “involves leadership responsibilities” and “work involves travelling,” while the latter is driven by “central location,” which is more frequently detected in our text analysis. [^23] See Appendix Table A.4, which shows the results of the same validation exercise for each of the 47 underlying the ten broad categories in Table 3. Overall, high levels of precision indicate that the attributes recovered by text analysis reflect the actual content of the attributes well. Sensitivity is always equal to or larger than precision, above 80 percent for all attributes.
3.3 The Prevalence of Advertised Job Attributes
We now use the detailed advertised job attributes to highlight several salient descriptive statistics. Figure 2 reports the share of job ads posted between 2015 and 2019 advertising each of the 47 attributes. About 55 percent of job ads provide some compensation-related information. However, explicit information about the actual level of pay is scarce. Less that 10 percent of ads mention a salary number or bracket (“compensation level”) and slightly more than 25 percent of ads indicate that pay is set by a collective wage bargaining agreement (“pay based on collective agreement”). [^24] By comparison, Batra et al. (2023) find that 13 percent of job ads have salary information in US data. Because these two attributes overlap to a large extent, 27 percent of job ads have information on the actual pay level of the position.
Figure 2: The Prevalence of Job Attributes Advertised in Vacancy Texts.
Note: This figure documents the prevalence of job attributes detected in job ads posted in Norway between 2015 and 2019 [N=863,478]. The light blue bars show the share of ads detected having each of the distinct job attributes. The dark blue bars show the share of ads detected with at least one attribute within ten broad categories. Appendix Figure A.3 shows the corresponding results for the 2002-2019 period as a whole.
Figure 3: Explained Variation in Publicly Advertised Job Attributes.
Note: This figure shows pseudo $R^{2}$ (McFadden, 1974) from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2015 and 2019 [N=863,478] on fixed effects denoting unique combinations of 2-digit industries (90 groups), location indicators (10 groups), 2-digit occupations (43 groups), and 108,730 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The last set of regressions controls for 563,170 unique combinations. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo. Appendix Figure A.4 shows the corresponding results for the 2002-2019 period as a whole.
Beyond information on the more tangible attributes of the job, such as pay, hours of work and contract duration, Figure 2 also shows that employers frequently choose to advertise more subjective aspects of the working environment, such as characteristics of the workplace (“workplace attributes”) or some appreciation of the type of tasks involved (“task-related attributes”). Around 30 percent of job ads describe these tasks as “challenging” or “interesting.” Similarly, 45 percent of vacancies feature some language about the quality of the work environment and more than 20 percent advertise “good colleagues.”
Next, we analyze to what extent differences in advertised attributes are explained by additional observable characteristics of the job ad, such as the industry of the posting establishment or the occupation associated with the job title stated in the job ad. In order to describe the variation in advertised job attributes that can be explained by such observable characteristics, we estimate a series of logistic regressions for the probability that a given attribute is advertised in the text of a vacancy with several sets of fixed effects: industries, locations, occupations, and establishments. Figure 3 reports the estimated pseudo-$R^{2}$, a summary measure of goodness of fit for binary outcomes (McFadden, 1974), from logistic regressions estimated separately for each attribute. [^25] As an alternative, we also estimated linear probability models using OLS with the same set of fixed effects separately for each attribute (results available upon request). However, as all of our pay and non-job attributes are binary, one could be concerned about interpreting standard goodness of fit from such regressions, which put restrictions on the range of the coefficient of determination $R^{2}$ (see, e.g., (Cox and Wermuth, 1992)). This measure therefore captures the variation in the content of job ads within industries, locations, occupations, and establishments.
A possibility is that the attributes advertised for a given position are largely determined by its industry or occupation, so the pay and non-pay attributes mentioned in vacancy texts directly reflect the type of job advertised. For example, it can be expected that working as a hospital nurse involves shift-work, in which case there is no value-added to mentioning this information in the corresponding job description. The data instead suggest that there is variation in advertised job attributes for very similar positions. Figure 3 shows that industries, occupations, and location fixed-effects explain at most 30 to 40 percent of the variation in the attributes advertised by employers across the 47 categories we retrieve. By contrast, when we add establishment fixed-effects, we find that the explained variation jumps up to 60 to 80 percent across attributes, which suggests that advertised attributes are in large part correlated across vacancies for establishments observed with several job ads. This last result is consistent with the existence of an “establishment fixed-effect” in advertised attributes, and it justifies the establishment-level analysis we develop in the next sections.
4 Estimating the Value of Employers
In this section, we estimate how workers value alternative employers using data on realized worker flows and wages. Our goal is to obtain measures of workers’ valuation of employers to which we can compare the information advertised in job ads. To this end, we estimate a version of the Burdett and Mortensen (1998) model in which employers post a value bundle using the revealed preference approach put forward by Sorkin (2018). In this setting, workers’ preferences over heterogeneous employers can be recovered from their mobility decisions. We expand on this methodology by adding information on the wages paid by employers, from which we can break down workers’ valuation of employers into a pay and non-pay component.
4.1 Model
Consider an economy in discrete time populated by a fixed measure of infinitely-lived workers and a fixed discrete number $J$ of employers. Workers have discount factor $\beta\in(0,1)$ and can be either employed at one of the $J$ employers or non-employed. Let $j\in\mathcal{J}=\{1,\dots,J\}$ denote employers and $V_{j}$ the value to a worker of being currently employed at employer $j$. Let $N$ denote non-employment and $V_{N}$ the value to a worker of being currently non-employed.
Both employed and non-employed workers search for better employers. Search is random. When searching, workers make contact with one of the $j\in\mathcal{J}$ employers with probability $f_{j}$, where $f$ captures the relative weight of employers in workers’ search. We maintain the wage-posting (here “value-posting”) assumption from Burdett and Mortensen (1998). The value of working at $j$ is summarized by $V_{j}$ and is not renegotiated even in the event of a change in the worker’s outside option.
Let $M_{jk}$ denote the measure of workers making a job-to-job move from employer $j$ to employer $k$. $M_{jk}$ is assumed to be the sum of involuntary relocation flows $M_{jk}^{R}$ and voluntary flows $M_{jk}^{V}$. Relocation job-to-job flows are given by
$M_{jk}^{R}=L_{j}\cdot\rho_{j}\cdot f_{k},$ (1)
where $L_{j}$ is employment at $j$, $\rho_{j}$ is the probability that workers employed at $j$ are forced to reallocate to another employer $j$, and $f_{k}$ is the probability that they find a job at employer $k$ [^26] There are no such relocation shocks in Burdett and Mortensen (1998), but they are a common feature in random search models with on-the-job search (see, e.g., Jolivet et al. (2006) and Bagger and Lentz (2019)). In Equation (1), workers do not make a choice and are forced to relocate to employer $k$. This is in contrast with voluntary job-to-job flows, which are given by
$M_{jk}^{V}=L_{j}\cdot(1-\rho_{j}-\delta_{j})\cdot\lambda_{1}\cdot f_{k}\cdot\text{Pr}(k\succ j).$ (2)
Voluntary worker flows from $j$ to $k$ are given by the measure of workers employed at $j$ ($L_{j}$), who are neither forced to reallocate (with probability $\rho_{j}$) nor forced to move to non-employment (with probability $\delta_{j}$), search for an alternative employer (with probability $\lambda_{1}$), and get an offer from employer $k$ with probability $f_{k}$.272727 The probabilities $\rho_{j}$ and $\delta_{j}$ are assumed to be mutually exclusive. In Equation (2), workers choose whether to remain with employer $j$ or move to $k$, which occurs with probability $\text{Pr}(k\succ j)$. This is the sense in which these flows are voluntary and reveal workers’ preferences over the set of employers $\mathcal{J}$. We follow Sorkin (2018) and assume there are i.i.d. taste shocks $\{\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$ that make the value of working at employer $j$ and of non-employment worker-specific in each period. We further assume that these shocks are drawn from a Gumbel distribution with location parameter normalized to zero and scale parameter $\sigma^{-1}$. [^28] Normalizing the location parameter to zero is without loss of generality because it shifts the value of working at each employer $j$ and of being in non-employment by the same amount. With this assumption, the acceptance probability for a worker with a job offer from employer $k$ currently working at $j$ is given by
$\text{Pr}(k\succ j)=\text{Pr}(V_{k}+\varepsilon_{k}\geq V_{j}+\varepsilon_{j})=\frac{\exp(\sigma V_{j})}{\exp(\sigma V_{j})+\exp(\sigma V_{k})}.$
The sequence of events described by the relocation and voluntary worker flows in Equations (1)-(2) implies the following expression for the value $V_{j}$ of working at employer $j$. Let $u_{j}$ denote the flow utility of working at $j$. $V_{j}$ can be expressed as
$\begin{split}V_{j}=u_{j}&+\beta\delta_{j}\mathbf{E}\big{[}V_{N}+\varepsilon_{N}\big{]}\\ &+\beta\rho_{j}\sum_{k\in\mathcal{J}}f_{k}\mathbf{E}\big{[}V_{k}+\varepsilon_{k}\big{]}\\ &+\beta(1-\delta_{j}-\rho_{j})\lambda_{1}\sum_{k\in\mathcal{J}}f_{k}\mathbf{E}\big{[}\max\big{\{}V_{j}+\varepsilon_{j},V_{k}+\varepsilon_{k}\big{\}}\big{]}\\ &+\beta(1-\delta_{j}-\rho_{j})(1-\lambda_{1})\mathbf{E}\big{[}\max\big{\{}V_{N}+\varepsilon_{N},V_{j}+\varepsilon_{j}\big{\}}\big{]}.\end{split}$ (3)
In Equation (3), the continuation value is made of the four following terms. With probability $\delta_{j}$, the worker transitions to non-employment, an exogenous shock. With probability $\rho_{j}$, the worker is hit by a relocation shock and is forced to move to an alternative employer by drawing from the offer distribution $f$, an exogenous shock. With probability $(1-\delta_{j}-\rho_{j})\lambda_{1}$, the worker gets an offer from a potential alternative employer by drawing from the offer distribution $f$ and decides whether to stay or move, an endogenous choice. With probability $(1-\delta_{j}-\rho_{j})(1-\lambda_{1})$, the worker decides whether to stay with their current employer or to move to non-employment, also an endogenous choice.
We depart from Sorkin (2018) by making an explicit functional form assumption on the utility flow $u_{j}$ of working at employer $j$ [^29] Sorkin (2018) makes a similar assumption on the value $V_{j}$ of employers; Lagos (2024) also follows a similar approach. By contrast, $V_{j}$ is not additively separable into a pay and non-pay in our formulation. We assume that $u_{j}$ depends linearly on a pay component $\ln W_{j}$ and a non-pay (or amenity) component $a_{j}$
$u_{j}=\ln W_{j}+a_{j}.$ (4)
The additive formulation $\ln W_{j}+a_{j}$ is standard in the literature that estimates static models of worker preferences over employers (Card et al., 2018; Lamadon et al., 2022). With an additional assumption to pin down the scale of the i.i.d taste shocks $\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$, we can recover the non-pay component $a_{j}$ associated with working at employer $j$ by inverting the value function in (3), conditional on the wage $W_{j}$ paid by employer $j$. We discuss this assumption below.
Given the form of the utility flow $u_{j}$ derived from working at employer $j$, workers’ valuation of $j$ can be summarized by the vector $\big{(}\ln W_{j},a_{j},\delta_{j},\rho_{j}\big{)}$. In what follows, we refer to $\ln W_{j}$ as the pay component, $a_{j}$ as the non-pay component, and $s_{j}=1-\delta_{j}-\rho_{j}$ as the job security component of workers’ preferences over employers.
We complete the description of the model by noting that there are similar expressions for workers flows to and from non-employment (counterparts to Equation 1 and Equation 2) and for the value of worker in non-employment (a counterpart to Equation 3), where workers in non-employment get a draw from the job offer distribution $\{f_{j}\}$ at rate $\lambda_{0}$. These expressions follow naturally and are relegated to Appendix C.
4.2 Identification
The model has several employer-level parameters, such as workers’ valuations of alternative employers $\{V_{j}\}_{j\in\mathcal{J}}$, and aggregate-level parameters, such as the offer arrival rate for employed workers $\lambda_{1}$. These parameters are identified using the matched employer-employee data described in Section 2, which contain the relevant information on wages and employer-to-employer flows. We now lay out the identification argument in a series of heuristic steps. At each step, we highlight the key assumptions required for identification.
Step 1: Parameters identified from worker flows
The employer-level parameters $\{\sigma V_{j},L_{j},f_{j},\rho_{j},\delta_{j}\}_{j\in\mathcal{J}}$ and the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are identified from (i) the realized mobility flows between employers $M_{jk}$, and (ii) which employers are shrinking or growing (see Sorkin, 2018, Section V.B). The notation $\sigma V_{j}$ emphasizes that worker flows pin down the value of working at employer $j$ up to the scale of the idiosyncratic taste shocks, which is identified in the next steps.
We briefly outline the main source of identification for each parameter intuitively and refer to the original paper for details. The main source of identification for $\sigma V_{j}$ is the aggregation of the realized mobility flows $M_{jk}$. An estimate for $L_{j}$ is obtained directly from the data as the share of worker-year observations at employer $j$. $f_{j}$ is identified from the share of non-employment flows to employer $j$. The separation shocks $\delta_{j}$ and $\rho_{j}$ are identified from the separation flows at shrinking firms, which give information on involuntary worker flows. Finally, the transition parameters $\lambda_{0}$ and $\lambda_{1}$ are pinned down by accounting for the aggregate transition flows, respectively from non-employment to employment and between employers.
Two sample restrictions are required for identification at this step. First, the aggregation of mobility flows only identifies $V_{j}$ within the strongly connected sets of employers.303030 Within a strongly connected set of employers, each employer has at least one worker moving in and one worker moving out. This condition is required for the fixed-point associated with the appropriately scaled matrix of worker flows to exist. See, e.g., Jackson (2010) for details on social network definitions. Second, each employer must hire at least one worker from non-employment for its sampling weight $f_{j}$ to be well-defined.
Step 2: Flow utility parameters
Given $\{\sigma V_{j},L_{j},f_{j},\rho_{j},\delta_{j}\}_{j\in\mathcal{J}}$ and the aggregate transition rate $\lambda_{1}$, $\sigma u_{j}$ can be recovered directly from Equation (3), again up to the scale of the i.i.d. taste shocks. This calculation is straightforward since, given that $\{\sigma\varepsilon_{i}\}_{i\in\{\mathcal{J},N\}}$ are i.i.d. draws from a Type 1 Extreme Value distribution with scale one, all expectations in Equation (3) admit the following closed-form solutions
$\begin{split}\sigma\mathbf{E}\big{[}V_{j}+\varepsilon_{j}\big{]}&=\sigma V_{j}+\gamma,\\ \sigma\mathbf{E}\big{[}\max\big{\{}V_{j}+\varepsilon_{j},V_{k}+\varepsilon_{k}\big{\}}\big{]}&=\ln\big{(}\exp(\sigma V_{j})+\exp(\sigma V_{k})\big{)}+\gamma,\end{split}$ (5)
where $\gamma$ denotes Euler’s constant. [^31] Gyetvai et al. (2022) make a related point in the context of a continuous time random search model. This step further requires to assume a value for the discount factor $\beta$, which we set in line with a 5 percent annual discount rate.
Step 3: Pay parameters
We make an additional set of assumptions on wage determination in the model such that the pay parameters $\ln(W_{j})$ are identified by the employer fixed-effect $\psi_{j(i,t)}$ in the standard two-sided unobserved heterogeneity regression
$\ln W_{it}=X_{it}^{\prime}\beta+\alpha_{i}+\psi_{j(i,t)}+e_{it},$ (6)
where $X_{it}$ are exogenous worker characteristics (such as age), $\alpha_{i}$ is the unobserved worker effect, $j(i,t)$ is worker $i$’s employer in period $t$, and $e_{it}$ is an idiosyncratic shock. $\psi_{j(i,t)}$ is only identified within the connected sets of employers, which is already required to identify $V_{j}$ from worker flows. [^32] A strongly connected set is connected by definition.
Specifically, we make the assumption that employers post piece-rate wage contracts as part of the utility bundle $V_{j}$. Piece-rate contracts are a commonly used assumption in random search models with wage-posting. [^33] See, for instance, Barlevy (2008) and Engbom and Moser (2022). This assumption implies that the wage premium paid by employers is log-additive in workers’ experience ($X_{it}$), ability ($\alpha_{i}$), and a time-varying and non-persistent productivity shock ($e_{it}$). Equation (6) is then correctly specified as long as individual-specific productivity does not interact with mobility decisions. We therefore also require that (i) individual-specific productivity is log-linear in the flow value of non-employment, and (ii) time-varying productivity shocks $e_{it}$ are mean-independent conditional on the realization of i.i.d. taste shocks. These assumptions ensure that individual-specific productivity does not affect the choice to move between employers and between employment and non-employment, in line with the model of mobility described so far.
Step 4: Non-pay parameters
Conditional on $u_{j}$ and a measure for $\ln W_{j}$, the non-pay parameters can be recovered by inverting Equation (4), which requires to pin down the parameter $\sigma$ controlling the scale of the i.i.d. taste shocks to recover $u_{j}$. We do this by assuming that the moment condition $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$ holds, so the variance in flow utility is the same as the variance in the flow utility of pay $\ln W_{j}$. An intuitive justification for this moment condition is that the flow utility of pay can be expected to be an important component of the flow utility of a job, so their variance should be of similar magnitude. However, this moment condition is not without loss of generality because it restricts the correlation between $\ln W_{j}$ and $a_{j}$ to be non-positive. It is immediate to check that $\text{Var}(u_{j})=\text{Var}\big{(}\ln W_{j}\big{)}$ implies $\text{Corr}(\ln W_{j},a_{j})\leq 0$. In this specific sense, the non-pay parameters we recover can be interpreted as having a compensating differentials component by definition.
4.3 Estimation
The model is estimated on the Norwegian matched employer-employee data described in Section 2. We use the establishment identifiers in the data as the direct counterpart to employer heterogeneity $j$ in the model. We follow the employer-employee fixed-effect literature and assign each individual one main employer in each year by selecting the establishment with their largest annual earnings. Our data feature the exact start date and end date of each employment spell with an employer in each year, so we can make a precise distinction between employer-to-employer moves (a data counterpart to $M_{jk}$ for moves between $j$ and $k$) and moves with a non-employment spell in-between (data counterparts to $M_{jN}$ and $M_{Nk}$ for moves between $j$ and $k$). Moves between two main employers that include an intervening spell with another employer also qualify as employer-to-employer moves. [^34] We define employer-to-employer flows as any two consecutive employment spells with a gap of at most 31 days.
Our sample restrictions follow from the model structure and the conditions required for identification. We restrict the sample to the period 2015-2019 and retain workers aged 20 to 60 (inclusive). We also restrict attention to relatively larger and more stable employers. Specifically, we require each establishment to be observed in at least two separate years and to have at least five non-singleton workers on average in each year, where a non-singleton worker is a worker observed at some later point within the sample period. In addition, the identification conditions outlined in Section 4.2 impose extra restrictions on the set of establishments included in the estimation sample. We restrict the sample to the largest strongly connected set of establishments and impose that each of these establishments hire at least one individual from non-employment. [^35] Focusing on the largest connected set is common in the literature. See, among others, Card et al. (2013) and Bonhomme et al. (2020) in the context of the two-sided unobserved heterogeneity regression model (6) and Sorkin (2018) and Morchio and Moser (2024) in the context of the fixed-point revealed preferences model. These last two restrictions are interdependent, and they are imposed recursively until the set of establishments converges.
In the model, we make the assumption that workers have identical preferences over employers and that they face similar search frictions. The piece-rate wage contract assumption allows wages to depend on a worker’s ability and experience, but the search parameters (such as the rate of arrival for employed workers $\lambda_{1}$ and offer distribution $\{f_{j}\}$) and preferences over employers are the same for all workers. As a result, there is no sorting on unobserved ability. Any two workers presented with the same choice of employer $j$ and $k$ make the same decision, up to their draw of taste shocks $(\varepsilon_{j},\varepsilon_{k})$. We allow for some degree of heterogeneity in preferences over employers and in search frictions by splitting the sample along observable pre-determined characteristics. We consider two demographic partitions of the data: women with and without a post-secondary education degree, and men with and without a post-secondary education degree. When splitting the sample by gender and by education, we impose the identification restrictions separately in each partition of the data.
There is a trade-off between allowing for a greater degree of observable worker heterogeneity and the mobility restrictions required for identification. Constant parameters and identical worker preferences call for considering shorter time spans and narrower worker groups (such as workers with a specific degree). Conversely, the requirement to focus on stable employers with sufficient mobility calls for longer samples and broader worker groups. In what follows, we consider the pooled sample as our baseline, but we also report results by gender and by education.
The model has employer-specific parameters $\{V_{j},L_{j},f_{j},\rho_{j},\delta_{j},\psi_{j},a_{j}\}_{j\in\mathcal{J}}$. We reduce the number of these parameters by grouping employers using a clustering algorithm. The goal is to increase the number of movers underlying each of these parameters to estimate them more accurately. We use a k-means algorithm to create $G$ employer groups from the $J$ employers in each selected sample in a first pre-estimation step (Bonhomme et al., 2019). The k-means algorithm requires two inputs: a vector of establishment-level variables on which the classification operates and a number of groups $G\leq J$. The vector of clustering variables includes wages (we use the empirical cumulative distribution function of wages in each establishment, computed at the deciles of the distribution of wages in the whole sample), job flows (we use the job creation and destruction flow rates, separately for direct transitions between employers and to and from non-employment), and some additional establishment characteristics (we use a set of indicators for the industry and occupation composition of the establishment, as well as its location). [^36] For location, we cut the distribution of municipality employment into deciles and construct indicators for the establishment’s municipality. For example, Oslo accounts for more than 10 percent of Norwegian employment, so there is a specific indicator for establishments in Oslo. We choose the number of clusters $G$ with the objective to allow for a large degree of employer heterogeneity while still significantly reducing the number of parameters to be estimated and set $G=\text{round}(J/50)$ in our baseline. [^37] Our baseline model thus features approximately $G=\text{round}(78,133/50)\approx 1,560$ unique employer clusters. We provide robustness of our findings using $G=\text{round}(J/25)$ and $G=\text{round}(J/100)$, respectively.
We obtain estimates of the establishment wage premium from the two-sided unobserved heterogeneity regression (6). We estimate this equation separately in each sample. Our wage measure is log average earnings per hour, where earnings are total pre-tax earnings and hours are total contract hours over the duration of each yearly employment spell. The vector of worker-level characteristics $X_{it}^{\prime}$ includes year-by-education fixed effects and a separate cubic age polynomial for each education group, restricted to be flat at age 40.383838These are the same set of worker characteristics as in Card et al. (2013). In line with our clustering strategy, we estimate wage premiums at the group level using the group identifiers obtained from the k-means algorithm.
As shown in Table 1, a minority of employers satisfy the stability and mobility requirements in our sample. Across samples, these conditions leave us with around 25 percent of the initial number of employers. But because of the long tail in the employer-size distribution, our estimation sample still retains 80 percent of the initial number of workers and worker-year observations.
4.4 Results
Figure 4 shows two-way scatter plots of the parameter estimates in the baseline model. Starting with the joint distribution of the estimated overall employer values $V_{j}$ and employer pay premiums $\psi_{j}$ in Panel A, the key object of interest in Sorkin (2018), we find that the correlation between these model components is 0.36.393939Sorkin (2018) finds a correlation of 0.53 between overall values $V_{j}$ and pay premiums $\psi_{j}$ based on US matched employer-employee data (see Table II in his paper). Besides a different institutional setting and time period, US data on job-to-job transitions is available only at quarterly frequency and lacks information on hours worked, while we use the exact start and end date of each spell and hourly wages. Next, Panel B shows the estimated overall employer values $V_{j}$ along the horizontal axis and the corresponding employer non-pay values $a_{j}$ along the vertical axis. We also find a positive association between $V_{j}$ and employer non-pay values $a_{j}$, especially at the lower end of the value distribution. By comparison, differences in pay premiums are the only source of employer heterogeneity in the canonical wage-posting model Burdett and Mortensen (1998), which implies $a_{j}=0$ at all employers, representing a useful benchmark to compare our results. In Panel C, we find a strong positive correlation of 0.68 between $V_{j}$ and the flow utility $u_{j}$ associated with being employed at $j$, capturing both pay and non-pay values, as shown in equation (4). In Panel D, we show the joint distribution of $V_{j}$ and job security value $s_{j}=1-\delta_{j}-\rho_{j}$, finding a strong positive correlation of 0.58. By contrast, in Burdett and Mortensen (1998), this correlation is also zero by construction since job security is assumed to be the same across all employers.
Overall, panels A-D show that there is a positive gradient between the value of employers $V_{j}$ and its pay, non-pay, and job security components. The plots also suggest that these relationships are to some degree non-linear. Differences in non-pay and job security are mostly a characteristic of employers at the lower end of the $V_{j}$ distribution, while differences in pay are mostly a characteristic of employers at the upper end of the $V_{j}$ distribution.
Figure 4: Model Estimates: Employer Pay, Non-Pay and Job Security Values.
Note: The figure provides two-way scatter plots depicting our baseline model estimates for employers’ overall value ($V_{j}$), flow utility value ($u_{j}$), pay value ($\psi_{j}$), non-pay value ($a_{j}$), and job security value ($s_{j}$). The size of the scatter plot markers are proportional to the number of worker-years in each cluster. Local linear regressions and correlations are weighted by worker-years. Dashed horizontal lines refers to the standard Burdett and Mortensen (1998) model, i.e., no non-pay value and common job security across employers. The latter is calculated as the (weighted) average of job security ($s_{j}$) across employers.
Next, we focus on the joint distributions between pay premiums $\psi_{j}$ and the other sources of employer heterogeneity that determine overall employer values in our model. Importantly, in Panel E, we find a strong negative correlation of $-0.64$ between the estimated employer pay $\psi_{j}$ and non-pay $a_{j}$ values, reflecting sizable compensating differentials in our framework. By comparison, in the Burdett and Mortensen (1998) model, there are no systematic non-pay differences across employers ($\text{Corr}(\psi_{j},a_{j})=0$). Since the correlation is clearly less than zero, there is still substantial variation in non-pay $a_{j}$ at every estimated level of employer pay $\psi_{j}$.
Further, in Panel G, we show that the estimated job security component $s_{j}$ and pay component $\psi_{j}$ are positively related, with a correlation of 0.3.404040This finding is in line with the estimates in Jarosch (2023), who estimates a positive correlation between match productivity and job security in a related random search framework with job-to-job transitions. In the canonical model, this correlation is also zero by assumption since job security is the same at all employers. Finally, we find that employer flow utilities $u_{j}$ are positively correlated with pay and non-pay values in Panels F and H, respectively, consistent with the findings in Panels A-C, while there does not appear to be a clear association between employer job security $s_{j}$ and employer non-pay and flow utility as shown in Panels I and J.
Table 4: Summary of Model Estimates.
Correlation between Employer Pay Value $\psi_{j}$ and:
Overall Value Job Security Value Non-Pay Value
$V_{j}$ $s_{j}$ $a_{j}$
(1) (2) (3)
Panel A. Baseline Model: $G=\text{round}(J/50)$
   Pooled sample 0.356 0.297 -0.639
   Women 0.193 0.307 -0.712
   Men 0.469 0.256 -0.579
   College 0.317 0.301 -0.634
   Non-college 0.333 0.299 -0.654
Panel B. Smaller Clusters: $G=\text{round}(J/25)$
   Pooled sample 0.334 0.255 -0.641
   Women 0.183 0.276 -0.708
   Men 0.430 0.204 -0.590
   College 0.296 0.252 -0.615
   Non-college 0.303 0.235 -0.630
Panel C. Larger Clusters: $G=\text{round}(J/100)$
   Pooled sample 0.371 0.323 -0.635
   Women 0.190 0.327 -0.721
   Men 0.500 0.299 -0.560
   College 0.343 0.358 -0.636
   Non-college 0.346 0.302 -0.645
Notes: The table shows key summary statistics of the estimated model parameters, namely the correlation between employer pay value and overall value (Corr($\psi_{j}$,$V_{j}$)), job security value (Corr($\psi_{j}$,$s_{j}$)), and non-pay value (Corr($\psi_{j}$,$a_{j}$)), respectively, across alternative specifications and sample restrictions. The model estimates are for the period 2015-2019.
In describing the model estimates by gender and by education, we focus on three of the main bi-variate relationships described in Figure 4: Corr($\psi_{j}$,$V_{j}$), Corr($\psi_{j}$,$s_{j}$) and Corr($\psi_{j}$,$a_{j}$). These correlations represent a good summary of the key quantitative properties of the model and are shown in Table 4. The first row in Panel A gives these correlations for the pooled sample, which correspond to the parameter estimates shown in Figure 4. Next, we report these measures by gender groups and education groups, based on separate estimations of the model for each group. The main differences are found across gender; we find a weaker correlation between employer pay and overall value for women, and consequently stronger compensating differentials for women as reflected by a more negative correlation between pay and non-pay values. Further, in Table 4, Panels B-C, we report the same correlations for alternative model specifications, where we change the number of clusters used to group employers. The overall patterns are remarkably similar when we set the number of clusters to round($J/25$) or round($J/100$), as opposed to round($J/50$) as in our baseline estimation.
To sum up, through the lens of a model with rich heterogeneity on the employer side, workers’ preferences over alternative employers are only partially captured by pay premiums. This conclusion is robust to several alternative specifications, as shown in Table 4. As a result, the content of job ads can be relevant for other dimensions of employer heterogeneity besides differences in pay within our framework, as we study next.
5 The Information Content of Job Ads
In this section, we bring together the job attributes advertised by employers extracted from vacancy texts in Section 3 and the value of employers estimated from worker mobility and wages in Section 4. We start this analysis by providing evidence on the job attributes that are publicly advertised by employers depending on their pay and non-pay values. We then analyze the predictive power of publicly advertised job attributes for employers’ actual values. Lastly, we provide a model-based measure of the information content of job ads.
5.1 What do Attractive Employers Offer in Job Ads?
Are attractive employers more likely to mention certain job attributes in publicly posted vacancies? We first investigate whether employers estimated as high-pay or low-pay advertise different types of attributes in their job postings. We estimate the regression model:
$\text{Attribute}_{g(j)}^{k}=\alpha_{0}^{k}+\alpha_{\psi}^{k}\cdot\psi_{g(j)}+\alpha_{X}^{\prime}X_{g(j)}+\varepsilon_{g(j)}^{k},$ (7)
at the establishment-cluster level $g$, where $\text{Attribute}_{g(j)}^{k}$ is the share of job ads from establishments in cluster $g$ that advertise attribute $k$, $\psi_{g(j)}$ is their estimated pay component, $X_{g(j)}$ is a vector of occupation, industry, and location controls, and $\varepsilon_{g(j)}^{k}$ is an error term. [^41] For each establishment-cluster g, the vector $X_{g(j)}$ controls flexibly for the composition of workers and establishments in our matched employer-employee data, with respect to 2-digit industries, 2-digit occupations, and 10 location groups. The patterns we describe below are robust to dropping controls $X_{g(j)}$. The coefficient of interest is $\alpha_{\psi}^{k}$, which measures to what extent employers with a standard deviation higher pay premium tend to advertise job attribute $k$ conditional on the controls in $X_{g(j)}$. We weight by the number of worker-years in each establishment-cluster.
We proceed similarly to study which attributes employers estimated as having high amenity value are more likely to advertise, substituting $a_{g(j)}$ for $\psi_{g(j)}$ in Equation (7), from which we obtain a second set of coefficients of interest $\alpha_{a}^{k}$ for each advertised job attribute $k$. While the pay and amenity estimates are negatively correlated, as shown in Panel C of Figure 4, there is substantial variation in the employer non-pay values also conditional on the employer pay premiums, and therefore, $\alpha_{a}^{k}$ is not proportional to $-\alpha_{\psi}^{k}$.
Figure 5: Publicly Advertised Job Attributes By Employer Pay and Non-Pay Values.
(a) Pay ($\psi_{j}$)
(b) Non-Pay ($a_{j}$)
Note: This figure shows parameter estimates from separate regressions of the share of each ad attribute on employer pay ($\psi_{j}$) and non-pay ($a_{j}$) values, capturing the change in the fraction of ads with each job attribute associated with a standard deviation increase in employer value. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group). Appendix Figure A.5 shows corresponding results by overall employer values ($V_{j}$).
Figure 5 reports the estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ for all attributes we derive from the vacancy data. Panel (a) shows that high-pay employers are more likely to mention characteristics associated with a more generous compensation package (“pension/insurance scheme,” “competitive pay,” “collective agreement pay”) and career opportunities (“career paths,” “on-the-job training”). Perhaps more surprisingly, we also find that several attributes related to the workplace, such as “good colleagues” and “social environment,” are associated with high-pay employers. At the other end of the spectrum, we see that most low-pay employers are more likely to specify the type of hours and the type of contract in their job postings. Panel (b) further shows that high non-pay employers are more likely to specify working hours and contract duration in the text of their job vacancies, especially in relation to a predictable schedule (“regular daytime schedule”) or flexible schedule (“possibility to work flexible hours”). Conversely, several career-related (“on-the-job training”) and workplace-related (“social environment,” “good colleagues”) attributes are negatively associated with non-pay values. Non-standard hours (“shift work”) is also more frequently advertised by employer with low non-pay value. Finally, Figure 5 shows that many estimated $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ are not statistically different from zero. We therefore fail to detect any correlation between many advertised attributes and the estimated value of the posting employers.
Overall, the correlations that we document in Figure 5 are consistent with the existence of compensating differentials, in line with the negative correlation between the pay and non-pay model estimates (Panel E in Figure 4). Notably, jobs offered by employers with higher non-pay value appear to more frequently have convenient hours, echoing the findings obtained in experimental settings (Mas and Pallais, 2017). However, the more accurate interpretation of the coefficients $\alpha_{\psi}^{k}$ and $\alpha_{a}^{k}$ is as measures of what high-pay and high non-pay employers are more likely to advertise in their ads. The advertised attributes need not translate directly into attributes of the actual jobs for the following reasons. First, it is unclear to what extent employers can deliver on many of these attributes, especially given that some of them are vague (“social environment,” “good colleagues”). Second, employers can omit specific characteristics from the text of the ad even though they are actual attributes of the position. As an example, they may actually offer a competitive salary without mentioning it explicitly in the corresponding vacancy text. As such, we are cautious in interpreting the pay and non-pay attributes mentioned in job ads as reflecting actual job attributes in general.
5.2 Predicting the Value of Employers from Job Ads
How predictive are publicly advertised job attributes of the employers’ actual pay and non-pay values? To assess the predictive power of the pay and non-pay content of job ads, we now estimate the following regression models, with the overall value of employers $V_{g(j)}$ or one of its components (i.e., pay $\psi_{g(j)}$ or non-pay $a_{g(j)}$) as the outcome variable:
$V_{g(j)}=\beta_{0}+\sum_{k}\beta_{k}\cdot\text{Attribute}_{g(j)}^{k}+\beta_{X}^{\prime}X_{g(j)}+\varepsilon_{g(j)}^{V}.$ (8)
In Equation (8), $\text{Attribute}_{g(j)}^{k}$ is the share of ads posted by employers in establishment-cluster $g$ that features job attribute $k$, $X_{g(j)}$ is a set of industry, location and occupation controls that is meant to capture other commonly observed characteristics listed in job ads, and $\varepsilon_{g(j)}^{V}$ is an error term. We weight by the number of worker-years in each cluster.
Figure 6: Predictive Power of Publicly Advertised Job Attributes.
Note: This figure shows adjusted $R^{2}$s from regressions of estimated model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two-digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). Appendix Figure A.7 provides corresponding results separately for each of the four five-year periods between 2000 and 2019.
Summarizing the overall predictive power job ads, Figure 6 shows the adjusted-$R^{2}$ of the regression model (8) with and without controlling for the variables in $X_{g}.$ [^42] The formula used to calculate predictive power is as follows: Adjusted $\textnormal{Adjusted }R^{2}=1-\frac{SS_{res}/df_{res}}{SS_{total}/df_{total}}$ where $SS_{res}$ is residual sum of squares, $df_{res}$ is the number of establishment clusters (approximately 1,560 in our baseline) minus the number of explanatory variables, and $SS_{total}$ and $df_{total}$ are from regressions on a constant term only. The partial $R^{2}$ is equal to the share of the variation that is unexplained in a regression on baseline controls only but explained when data extracted from job ads is added to the right hand side: Partial $\textnormal{Partial }R^{2}=\frac{R^{2}_{full}-R^{2}_{baseline}}{1-R^{2}_{baseline}}$ where $R^{2}_{full}$ ($R^{2}_{baseline}$) is adjusted-$R^{2}$ from a regression on baseline controls and (without) job attributes. This plot shows that the information contained in job ads meaningfully predicts overall employer values and their components. We find that the non-pay attributes (blue bars) are about as predictive of the overall employer values as the pay-related attributes (red bars). In total, solely based on the job attributes we are able to detect in ads (i.e., without including the controls in $X_{g(j)}$), we are able to explain about 60 percent of the variation in employer pay and their overall values, and about 50 percent of non-pay and 40 percent of the variation in job security. Once we purge the information contained in $X_{g}$, we find that detected job attributes still explain around 10 to 20 percent of the remaining variation in employer values, as shown by the partial-$R^{2}$ in Figure 6.
We further decompose the adjusted-$R^{2}$ measures reported above into contributions from different categories of job attributes in Appendix Figure A.6. Panel (a) documents that financial attributes explain a large share of the variation in employer pay values, while hours of work and convenient hours contain some of the residual variation. Similarly, Panel (b) shows that a large fraction of the variation in non-pay employer values can be explained by detected job attributes related to convenient and inconvenient hours, while pay-related attributes contribute less to explanatory power. In a similar spirit, in Appendix D, we further document that the patterns discussed here are robust to focusing on specific job attributes and the order in which we add various attributes. Specifically, we show evidence from a series of regressions where we sequentially add the individual job attributes that provide the largest increase in predictive power. Based on this approach, we confirm that a substantial share of the predictive power with respect to employer pay values is contained in the information related to financial attributes, while the information on convenient or inconvenient hours remains key in the predictions of employer non-pay values.
Next, we consider how the posted employer values that are predicted using Equation (8) relate to the average duration of vacancies posted by the corresponding employers. Using data on job ads, recent evidence by Mueller et al. (2024) and Bassier et al. (2023) shows that employer pay premiums and posted pay are negatively related to the duration of posted vacancies, consistent with the notion that high-pay employers are more attractive in the labor market. In Appendix Table A.5, Column (1), we show that the average vacancy duration is 15.4 percent shorter for job ads with a standard deviation higher overall posted employer value, while controlling flexibly for industries, occupations and locations listed in the job ads. Decomposing the overall posted employer value in its various components, we find a vacancy duration elasticity of $-0.4$ with respect to the posted pay premium and an elasticity of $-0.5$ with respect to the posted amenity value. Further, we get a negative relationship between posted job security and vacancy duration, finding 8.6 percent shorter average duration for job ads with a standard deviation higher job security value. These results confirm that both pay premiums, amenity values and job security shape the overall attractiveness of employers.
Heterogeneity by Gender and by Education
Returning to the regression model in Equation (8), Figure 7 shows the adjusted-$R^{2}$ by gender and by education, with and without controlling for the variables in $X_{g(j)}$. The outcomes in these regressions–the employer values–are estimated separately for each sub-sample, and all regressions are weighted by the corresponding number of worker-years in each sub-sample. These plots show that the information contained in vacancy texts meaningfully predicts the value of employers in all sub-samples. Across sub-samples, the magnitude of the overall $R^{2}$s are broadly similar to those reported for the pooled sample in Figure 6, i.e., around 60 percent for the estimated overall employer values. The partial $R^{2}$s tend to be slightly larger when we split the sample by gender and by education. It is consistently greater than 20 percent for the overall employer values as compared to 15 percent in the pooled sample.
Figure 7: Heterogeneity in the Predictive Power of Publicly Advertised Job Attributes.
(a) Men
(b) Women
(c) College
(d) Non-College
Note: This figure shows adjusted $R^{2}$s from regressions of estimated population group-specific model parameters on pay and non-pay job attributes and other ad characteristics retrieved in the text analysis, including the number of words in the ad, indicators for ad posted online, disclosure of company name, and the overall number of attributes in the ad. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels (a)-(b)) and education (Panels (c)-(d)), and are weighted by the number of worker-years in each cluster and population group. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation, and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group). Appendix Figure A.8 shows a corresponding plot with weights fixed across this sample split using the overall number of worker-years.
Overall, Figure 7 suggests that the most salient difference across sub-samples in terms of the explanatory power of the content of job ads is found between college and non-college workers. In particular, the explanatory power of ads is at least 10 percentage points lower for the estimated overall and non-pay employer values for non-college workers relative to college workers, both for the overall and partial $R^{2}$. By comparison, the overall $R^{2}$ suggests that the content of job ads has a higher predictive power in terms of employers’ non-pay value for women than for men. However, this result disappears once we control for $X_{g(j)}$. In Appendix Table A.5, Columns (2)-(5), we report the elasticity of vacancy duration with respect to posted pay and amenity values for each sub-sample, finding broadly similar patterns.
5.3 A Model-Based Measure of Ad Informativeness
What is the information content of job ads in terms of pay and non-pay attributes offered by employers? We now use the link between the estimated value of employers and the information extracted from vacancy texts to provide a measure of ad informativeness based on the model introduced in Section 4. Within the model, the $R^{2}$ calculations can be translated into upward and downward mobility flows. Specifically, we introduce the notation
$\widehat{M}_{jk}=L_{j}\cdot\Big{[}\rho_{j}\cdot f_{k}+(1-\rho_{j}-\delta_{j})\cdot\lambda_{1}\cdot f_{k}\cdot\text{Pr}(\widehat{V}_{k}\geq V_{j})\Big{]},$ (9)
for the job-to-job flows from employer $j$ to employer $k$, where $\widehat{V}_{k}$ is the value of employers predicted from one of the regression models in Equation (8). [^43] The regression models (8) are weighted by the job offer distribution $f_{j}$. This ensures that the average job offer $\sum_{j^{\prime}}f_{j^{\prime}}\widehat{V}_{j^{\prime}}$ remains the same across specification by construction. In words, we interpret $\widehat{M}_{jk}$ as the counterfactual flows from employer $j$ to employer $k$ if workers were to use the information advertised on employer $k$’s vacancies to decide whether to move. We further make a distinction between the job-to-job flows where the actual value of the destination employer $V_{k}$ is greater or lower than the value of employer $j$, which we denote, respectively, $\widehat{M}_{jk}(\uparrow)=\widehat{M}_{jk}\cdot\mathbf{1}\{V_{k}\geq V_{j}\}$ and $\widehat{M}_{jk}(\downarrow)=\widehat{M}_{jk}\cdot\mathbf{1}\{V_{k}<V_{j}\}$.
We compute economy-wide summary measures of overall counterfactual job-to-job mobility by aggregating across all bilateral employer flows as $\widehat{M}=\sum_{j^{\prime},k^{\prime}}\widehat{M}_{j^{\prime}k^{\prime}}$ for each alternative prediction model. We proceed similarly to aggregate job-to-job flows to employers whose actual value is above and below the worker’s current employer, with, respectively, $\widehat{M}(\uparrow)=\sum_{j,k}\widehat{M}_{jk}(\uparrow)$ providing a summary measure of upward job mobility and $\widehat{M}(\downarrow)=\sum_{j,k}\widehat{M}_{jk}(\downarrow)$ providing a summary measure of downward job mobility.
Table 5: Counterfactual Analysis: Informativeness of Job Ads for Job-to-Job Mobility.
Relative Change in Job-to-Job Mobility (in %):
Pay Attributes Pay and Non-Pay Attributes Full Model
$\Delta$ Mobility $\Delta$ Mobility $\Delta$ Mobility
Overall $\uparrow$ $\downarrow$ Overall $\uparrow$ $\downarrow$ Overall $\uparrow$ $\downarrow$
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Change Relative to Benchmark Scenario:
Pooled sample 0.2 0.7 -0.5 0.6 1.6 -1.0 1.7 6.7 -5.0
Gender:
   Women 0.0 0.7 -0.7 0.3 1.8 -1.5 0.6 5.2 -4.6
   Men 0.3 0.4 -0.1 0.7 1.2 -0.5 2.1 5.8 -3.7
Education:
   College 0.1 0.8 -0.6 1.0 2.5 -1.5 2.2 7.0 -4.8
   Non-college 0.6 0.9 -0.3 1.0 1.6 -0.5 2.3 6.2 -3.9
Note: The table reports relative changes in job-to-job mobility rates (in percent) under alternative information sets used to predict the overall employer values, as compared to a benchmark scenario where workers can only use a linear projection based on industry, occupation, and location. In Columns (1)-(3), we report mobility rates where workers use pay information extracted from job ads, besides industry, occupation and location, while in Columns (4)-(6), workers use information on both pay and non-pay attributes as well as other ad characteristics. Finally, in Columns (7)-(9), we use the actual employer values obtained from the model. The columns denoted by $\uparrow$ and $\downarrow$ show the rates of changes in upward job mobility ($V_{k}$geq $V_{j}$) and downward job mobility ($V_{k}$ $<$ $V_{j}$), respectively.
In Table 5, we investigate how job-to-job mobility changes under alternative information sets used to predict overall employer values. All results are shown relative to the benchmark linear prediction model in Equation (8) with occupation, industry, and location controls, but no information from vacancy texts. The numbers reported in the top row suggest that having detailed information on job attributes from job ads would increase (reduce) inflows of workers to better (worse) jobs by 1.6 (1.0) percent, as compared to the benchmark scenario where only “one-liner” vacancies listing occupation, industry, and location are posted. This suggests that publicly advertised information on both pay and non-pay attributes can have important implications for the strength and direction of job-to-job flows. Using our full model, we find that upward mobility goes up by 6.7 percent relative to the benchmark. The corresponding figure for downward mobility is 5 percent. This result suggests that publicly advertised information on job attributes explains around 25 percent of the upward mobility in job-to-job flows that we observe in our model. Using this measure of ad informativeness, we also find evidence of heterogeneity by gender and by education. The heterogeneity results in Table 5 imply that the information on job attributes in ads explains more of the upward mobility in job-to-job flows for women than for men (30 percent vs 20 percent) and for college graduates than for non-college workers (35 percent vs 25 percent).
In Table 6, we extend the analysis in Table 5 to upward and downward mobility in employer pay and non-pay values. Using the same linear models to define counterfactual mobility flows, we now quantify to what extent detailed ad information translates into additional mobility toward better-paying employers $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{\psi_{k}\geq\psi_{j}\}$ or employers offering better amenities $\sum_{j,k}\widehat{M}_{jk}\cdot\mathbf{1}\{a_{k}\geq a_{j}\}$. We similarly define downward mobility in pay and non-pay.
Table 6: Counterfactual Analysis: Informativeness of Job Ads for Job-to-Job Mobility.
Relative Change in Job-to-Job Mobility (in %):
Pay Pay and Non-Pay
Attributes Attributes Full Model
$\Delta$ Mobility $\Delta$ Mobility $\Delta$ Mobility
$\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$
(1) (2) (3) (4) (5) (6)
Change Relative to Benchmark Scenario:
Pooled sample
   Pay 0.5 -0.2 0.4 0.3 0.7 0.1
   Non-pay -0.1 0.3 0.5 0.1 2.1 -0.4
Gender:
   Women
    Pay -0.3 0.3 -0.6 0.9 -0.2 0.8
    Non-pay 0.4 -0.4 1.0 -0.8 1.4 -0.8
   Men
    Pay 0.4 -0.2 0.5 0.2 1.1 1.1
    Non-pay 0.0 0.3 0.4 0.3 2.1 0.0
Education:
   College
    Pay 0.1 0.1 0.2 0.8 1.0 1.1
    Non-pay 0.4 -0.3 1.6 -0.6 2.7 -0.6
   Non-college
    Pay 1.3 -0.6 1.4 -0.4 1.8 0.5
    Non-pay -0.1 0.7 0.3 0.8 1.7 0.6
Note: The table reports relative changes in job-to-job mobility rates (in %) under alternative information sets used to predict the overall employer values, as compared to a benchmark scenario where workers can only use a linear projection based on industry, occupation and location. Each row reports upward and downward mobility in Pay ($\psi_{j}$) or Non-pay ($a_{j}$) for each sub-sample. The columns denoted by $\uparrow$ shows the rate of change in upward mobility in Pay ($\psi_{k}$geq $\psi_{j}$) and Non-pay ($a_{k}$geq $a_{j}$), respectively. The columns denoted by $\downarrow$ shows the rate of change in downward mobility in Pay ($\psi_{k}$ $<$ $\psi_{j}$) and Non-pay ($a_{k}$ $<$ $a_{j}$), respectively.
We draw two main conclusions from this exercise. First, we find that the magnitude of the changes in job-to-job flows to employers offering better pay and better non-pay is smaller across prediction models, as compared to Table 5. For instance, the increase in job-to-job flows to employers with higher value is 1.6 percent using the predictive model with both pay and non-pay attributes in the pooled sample (column (5) top row in Table 5). This number drops to 0.4 percent (0.5 percent) for flows to employers offering better pay (non-pay). This reflects the relatively limited correlation between the overall value of employers $V_{j}$ and its components: the correlation between $\psi_{j}$ and $V_{j}$ is 0.36 in the pooled sample (see Table 4). Second, this exercise suggests that there is heterogeneity in the direction of job-to-job flows predicted by the pay and non-pay attributes contained in vacancy texts. Focusing on gender differences, we find that men are more likely to move to employers offering better pay as their information set is expanded. Relative to the baseline model, their job-to-job mobility to better paying employers increases by 0.5 percent using the full set of attributes advertised in vacancies (Column (3) in Table 6). By contrast, women’s job-to-job mobility to employers offering higher pay decreases with the full set of advertised attributes. The flip side of this decrease in job-to-job flows to higher-paying employers is a one percent increase in the flows to employers offering better amenities. The intuition behind the gender results is straightforward. In the estimated model, pay is a more important component of the value of employers for men than for women, as shown in Table 4. We also find that advertised attributes are slightly more informative of the value of employers for women than for men (Table 5). As a result, with more information, women self-select more into high-amenity jobs. This last finding is consistent with gender differences in preferences over employer attributes carrying over to the job search stage (e.g., Le Barbanchon et al., 2021; Fluchtmann et al., 2024).
6 Conclusion
In this paper, we study the pay and non-pay content of job ads. Our first set of results is derived from analyzing the texts of a comprehensive database of job ads. We find that the texts of these ads specify many job attributes relevant to workers’ preferences over alternative employers. Almost every job ad mentions at least one of these attributes with an average of 6.6 attributes per ad. By considering a wide range of advertised job attributes, our analysis adds to prior work focusing on specific content, such as salary information (Batra et al., 2023), remote work (Hansen et al., 2023), or flexible working arrangements (Adams-Prassl et al., 2023a).
We also recover workers’ valuation for alternative employers using a structural model estimated on matched employer-employee data. Our second set of results follows from linking these estimates to the pay and non-pay content of job ads. High-pay employers are more likely to advertise aspects of the compensation package, while high-amenity employers more often mention working arrangements, such as contract duration and hours worked, in their postings. We confirm using several alternative metrics that the content of job ads has predictive power for workers’ valuation of employers above other observable characteristics, such as industries and occupations. Taken together, these results suggest that the pay and non-pay content of job vacancies represents one reliable source of information for workers assessing their options in the labor market.
Our approach based on extracting a wide range of job attributes from vacancy texts points to several directions for future work. First, the pay and non-pay content of job postings can be expected to differ across labor markets. As one example, health insurance is largely absent from job ads in Norway, where healthcare is universal, but it can be expected to be an important job attribute wherever employers are commonly involved, such as in France or the US. Second, there might be a cyclical element to the content of job ads, with recruiters potentially advertising more and different attributes in tight labor markets. Third, while prior experimental work has focused on the salary component of vacancies (Belot et al., 2022), our results show that there is scope to manipulate a much broader range of job attributes to further unpack the search behavior of job seekers.
Finally, our results have implications for the design of policies related to the information contained in job ads. Several jurisdictions have passed legislation mandating employers to disclose a salary range in their job openings. Our analysis suggests that there is scope to regulate the provision of additional job characteristics besides pay, such as information on working arrangements. Similarly, these results can be useful for the design of online job platforms in that they provide guidance on the attributes employers should be encouraged to provide in their postings.
Appendix
Appendix A Additional Tables and Figures
Table A.1: Overview of Sample Selection: Five-Year Periods, Cluster-Level Match.
Job Ads Establishments Workers Worker-Years
(1) (2) (3) (4)
Panel A: 2000-2004
All Observations $339,791$ (100%) $210,958$ (100%) $2,468,402$ (100%) $9,820,699$ (100%)
Observe Vacancy Text $330,374$ (97.2%) $87,790$ (41.6%) $2,077,549$ (84.2%) $8,488,411$ (86.4%)
Excluding Staffing Agencies $323,332$ (95.2%) $87,789$ (41.6%) $2,077,548$ (84.2%) $8,488,320$ (86.4%)
Strongly Connected Set $221,999$ (65.3%) $54,759$ (26.0%) $1,934,829$ (78.4%) $7,915,843$ (80.6%)
All Sample Restrictions $209,798$ (61.7%) $54,752$ (26.0%) $1,934,476$ (78.4%) $7,914,290$ (80.6%)
Panel B: 2005-2009
All Observations $948,589$ (100%) $228,287$ (100%) $2,663,696$ (100%) $10,706,270$ (100%)
Observe Vacancy Text $917,501$ (96.7%) $119,000$ (52.1%) $2,389,467$ (89.7%) $9,727,646$ (90.9%)
Excluding Staffing Agencies $850,809$ (89.7%) $119,000$ (52.1%) $2,389,467$ (89.7%) $9,727,646$ (90.9%)
Strongly Connected Set $708,646$ (74.7%) $61,459$ (26.9%) $2,134,368$ (80.1%) $8,802,142$ (82.2%)
All Sample Restrictions $631,441$ (66.6%) $61,453$ (26.9%) $2,134,333$ (80.1%) $8,801,945$ (82.2%)
Panel C: 2010-2014
All Observations $734,996$ (100%) $237,677$ (100%) $2,854,890$ (100%) $11,797,654$ (100%)
Observe Vacancy Text $702,292$ (95.6%) $110,132$ (46.3%) $2,513,019$ (88.0%) $10,570,006$ (89.6%)
Excluding Staffing Agencies $670,108$ (91.2%) $110,132$ (46.3%) $2,513,019$ (88.0%) $10,570,006$ (89.6%)
Strongly Connected Set $572,520$ (77.9%) $62,934$ (26.5%) $2,264,686$ (79.3%) $9,632,981$ (81.7%)
All Sample Restrictions $521,661$ (71.0%) $62,933$ (26.5%) $2,264,662$ (79.3%) $9,632,258$ (81.6%)
Panel D: 2015-2019
All Observations $915,790$ (100%) $301,716$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $127,913$ (42.4%) $2,733,050$ (87.5%) $11,898,501$ (89.6%)
Excluding Staffing Agencies $774,076$ (84.5%) $127,904$ (42.4%) $2,732,297$ (87.5%) $11,895,973$ (89.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $78,118$ (25.9%) $2,499,506$ (80.0%) $10,974,896$ (82.7%)
Panel E: 2000-2019
All Observations $2,939,166$ (100%) $472,945$ (100%) 3,867$,23$ 867 , 232 (100%) $45,602,960$ (100%)
Observe Vacancy Text $2,813,645$ (95.7%) $222,739$ (47.1%) $3,722,456$ (96.3%) $40,684,564$ (89.2%)
Excluding Staffing Agencies $2,618,325$ (89.1%) $222,739$ (47.1%) $3,722,316$ (96.3%) $40,681,944$ (89.2%)
Strongly Connected Set $2,247,273$ (76.5%) $126,925$ (26.8%) $3,621,634$ (93.6%) $37,328,744$ (81.9%)
All Sample Restrictions $1,987,892$ (67.6%) $126,914$ (26.8%) $3,621,402$ (93.6%) $37,323,388$ (81.8%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for each of the four five-year periods (Panels A-D) and the combined 2000-2019 period (Panel E). “All Observations” refers to observations with valid establishment identifiers in each five-year period. “Excluding Staffing Agencies” are observations linked to at least one vacancy with observed text not posted by a recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural model. The sample sizes are calculated after matching vacancies to the data on the establishment-cluster level in the corresponding five-year period. In Panel A, vacancy texts are only observed for 2002-2004 due to the availability of data on job ads, and similarly, in Panel E, vacancy texts are observed between 2002 and 2019.
Table A.2: Overview of Sample Selection: Five-Year Periods, Establishment-Level Match.
Job Ads Establishments Workers Worker-Years
(1) (2) (3) (4)
Panel A: 2000-2004
All Observations $339,791$ (100%) $210,957$ (100%) $2,468,402$ (100%) $9,820,699$ (100%)
Observe Vacancy Text $330,374$ (97.2%) $65,578$ (31.1%) $1,613,285$ (65.4%) $6,623,923$ (67.4%)
Excluding Staffing Agencies $323,332$ (95.2%) $65,264$ (30.9%) $1,589,271$ (64.4%) $6,553,017$ (66.7%)
Strongly Connected Set $221,999$ (65.3%) $54,759$ (26.0%) $1,934,829$ (78.4%) $7,915,843$ (80.6%)
All Sample Restrictions $209,798$ (61.7%) $32,342$ (15.3%) $1,446,473$ (58.6%) $5,980,817$ (60.9%)
Panel B: 2005-2009
All Observations $948,589$ (100%) $228,286$ (100%) $2,663,696$ (100%) $10,706,270$ (100%)
Observe Vacancy Text $917,501$ (96.7%) $105,358$ (46.2%) $2,165,790$ (81.3%) $8,822,007$ (82.4%)
Excluding Staffing Agencies $850,809$ (89.7%) $104,434$ (45.7%) $2,123,522$ (79.7%) $8,696,302$ (81.2%)
Strongly Connected Set $708,646$ (74.7%) $61,459$ (26.9%) $2,134,368$ (80.1%) $8,802,142$ (82.2%)
All Sample Restrictions $631,441$ (66.6%) $47,293$ (20.7%) $1,869,473$ (70.2%) $7,774,969$ (72.6%)
Panel C: 2010-2014
All Observations $734,996$ (100%) $237,676$ (100%) $2,854,890$ (100%) $11,797,654$ (100%)
Observe Vacancy Text $702,292$ (95.6%) $92,453$ (38.9%) $2,200,499$ (77.1%) $9,285,465$ (78.7%)
Excluding Staffing Agencies $670,108$ (91.2%) $91,377$ (38.4%) $2,146,382$ (75.2%) $9,113,678$ (77.2%)
Strongly Connected Set $572,520$ (77.9%) $62,934$ (26.5%) $2,264,686$ (79.3%) $9,632,981$ (81.7%)
All Sample Restrictions $521,661$ (71.0%) $44,661$ (18.8%) $1,899,569$ (66.5%) $8,180,234$ (69.3%)
Panel D: 2015-2019
All Observations $915,790$ (100%) $301,715$ (100%) $3,123,055$ (100%) $13,278,335$ (100%)
Observe Vacancy Text $863,478$ (94.3%) $108,730$ (36.0%) $2,456,581$ (78.7%) $10,770,665$ (81.1%)
Excluding Staffing Agencies $774,076$ (84.5%) $107,204$ (35.5%) $2,401,811$ (76.9%) $10,566,082$ (79.6%)
Strongly Connected Set $744,108$ (81.3%) $78,133$ (25.9%) $2,500,325$ (80.1%) $10,977,778$ (82.7%)
All Sample Restrictions $624,992$ (68.2%) $58,153$ (19.3%) $2,170,687$ (69.5%) $9,651,061$ (72.7%)
Panel E: 2000-2019
All Observations $2,939,166$ (100%) $472,945$ (100%) 3,867$,23$ 867 , 232 (100%) $45,602,960$ (100%)
Observe Vacancy Text $2,813,645$ (95.7%) $200,220$ (42.3%) $3,589,923$ (92.8%) $35,502,060$ (77.9%)
Excluding Staffing Agencies $2,618,325$ (89.1%) $197,986$ (41.9%) $3,567,109$ (92.2%) $34,929,080$ (76.6%)
Strongly Connected Set $2,247,273$ (76.5%) $126,925$ (26.8%) $3,621,634$ (93.6%) $37,328,744$ (81.9%)
All Sample Restrictions $1,987,892$ (67.6%) $94,530$ (20.0%) $3,419,517$ (88.4%) $31,587,080$ (69.3%)
Note: This table documents the sample sizes for different sample restrictions imposed in the estimation for each of the four five-year periods (Panels A-D) and the 2000-2019 period (Panel E). “All Observations” refers to observations with valid establishment identifiers in each five-year period. “Excluding Staffing Agencies” are observations linked to at least one vacancy with observed text not posted by a recruitment or temporary employment agency. “Strongly Connected Set” refers to the strongly connected set of employers used in the structural model. The sample sizes in the last row in each panel are calculated after matching vacancies to the data on the establishment level in the corresponding five-year period (i.e., this excludes all establishments without posted vacancies in the corresponding five-year period). In Panel A, vacancy texts are only observed for 2002-2004 due to the availability of data on job ads, and similarly, in Panel E, vacancy texts are observed between 2002 and 2019.
Table A.3: Overview of Establishment Characteristics.
All Establishments A Observe Vacancy Text Excluding Staffing Agencies Strongly Connected Set Both Sample Restrictions
Panel A: Worker Characteristics
Women $43.7$% $49.9$% $49.9$% $51.2$% $51.2$%
College $34.9$% $37.9$% $37.9$% $40.2$% $40.2$%
Hourly Wage $25.34$ $25.16$ $25.15$ $25.45$ $25.45$
Panel B: Establishment Characteristics
Size $13.2$ $22.5$ $22.6$ $30.1$ $30.1$
Public Sector $17.3$% $25.8$% $25.8$% $30.7$% $30.7$%
Panel C: Industry Composition
Agriculture, Forestry and Fishing $2.8$% $1.3$% $1.3$% $1.0$% $1.0$%
Mining and Quarrying $0.4$% $0.4$% $0.4$% $0.5$% $0.5$%
Manufacturing $5.8$% $6.4$% $6.5$% $6.6$% $6.6$%
Energy Supply $0.4$% $0.5$% $0.5$% $0.4$% $0.4$%
Water Supply $0.5$% $0.6$% $0.6$% $0.6$% $0.6$%
Construction $10.3$% $8.2$% $8.2$% $8.0$% $8.0$%
Wholesale and Retail Trade $23.3$% $24.6$% $24.6$% $22.9$% $22.9$%
Transportation and Storage $5.5$% $4.5$% $4.5$% $4.6$% $4.6$%
Accommodation and Food Service $4.2$% $4.9$% $4.9$% $4.7$% $4.7$%
Information and Communication $3.1$% $2.8$% $2.8$% $2.9$% $2.9$%
Financial and Insurance $1.4$% $1.5$% $1.5$% $1.5$% $1.5$%
Real Estate $3.4$% $1.5$% $1.5$% $1.1$% $1.1$%
Professional, Scientific and Technical $8.2$% $5.8$% $5.8$% $5.1$% $5.1$%
Administrative and Support Service $4.0$% $4.2$% $4.0$% $4.4$% $4.4$%
Public Administration and Defense $2.9$% $4.2$% $4.2$% $4.3$% $4.3$%
Education $3.8$% $5.7$% $5.7$% $7.1$% $7.1$%
Health and Social Work $12.9$% $17.3$% $17.4$% $20.0$% $20.0$%
Arts, Entertainment and Recreation $2.0$% $1.8$% $1.8$% $1.7$% $1.7$%
Other Service Activities $4.3$% $3.6$% $3.6$% $2.5$% $2.5$%
Other Industries $0.3$% $0.0$% $0.0$% $0.0$% $0.0$%
Note: This table documents how the sample restrictions imposed in the estimation affects the composition of establishments between 2000 and 2019. Each panel shows averages over establishment-years, meaning that employers that are present in all years are implicitly weighted more than employer that are present in half of the years. Panel A shows average worker characteristics across establishment-years, weighting workers in small employers more than large employers. The first row, for example, is constructed by first calculating the average share of women within employer-years, and then averaging over employer-years. Panel C documents industry compositions using Classification of Economic Activities in the European Community (NACE).
Table A.4: Validation of Detected Pay and Non-Pay Job Attributes.
All Ads Job Ads in the Validation Sample
Text Analysis Text Analysis Manual Recognition Success Rate Precision Rate Sensitivity Rate
(1) (2) (3) (4) (5) (6)
Compensation Scheme
   –– Competitive Pay 14.9 17.5 19.2 94.2 88.6 80.5
   –– Collective Agreement Pay 23.3 31.5 33.8 87.8 84.1 78.5
   –– Incentive Pay Scheme 2.6 2.0 1.5 98.5 50.0 66.7
   –– Hiring Bonus 0.2 0.0 0.0 100.0
   –– Good Overtime Pay 0.4 0.2 0.8 99.5 100.0 33.3
   –– Any Other Mention of Pay 14.8 15.0 8.8 85.8 31.7 54.3
Financial Attributes
   –– Pension Scheme 28.6 34.8 32.8 98.0 94.2 100.0
   –– Insurance Scheme 18.0 23.8 21.5 96.2 87.4 96.5
   –– Mortgage Possibility 1.3 2.0 3.8 98.2 100.0 53.3
Career Opportunities
   –– Good Career Paths 11.9 16.8 6.2 85.0 23.9 64.0
   –– On-the-Job Training 39.2 46.8 39.2 78.0 68.4 81.5
Hours of Work
   –– Full-time Contract 58.8 55.8 53.2 95.0 93.3 97.7
   –– Part-time Contract 30.0 34.0 28.2 91.2 78.7 94.7
   –– Full-time/Part-time Choice 4.0 3.0 3.0 94.0 0.0 0.0
   –– On-call Employment 3.3 3.8 2.2 98.0 53.3 88.9
Convenient Hours
   –– Possibility to Work Flexible Hours 6.1 6.8 7.5 98.2 92.6 83.3
   –– Regular Daytime Work Schedule 14.2 8.8 8.8 97.0 82.9 82.9
   –– Exempt from Work Hour Regulations 0.4 1.2 0.2 98.5 0.0 0.0
Inconvenient Hours
   –– Shift Work 14.7 15.0 11.8 95.8 75.0 95.7
   –– Weekend/Evening/Night Work 17.4 21.8 17.0 94.8 77.0 98.5
   –– Overtime Work Required 0.2 0.0 0.2 99.8 0.0
Contract Duration
   –– Permanent Job 39.5 38.8 40.8 93.0 93.5 89.0
   –– Temporary Job 29.0 32.2 29.2 91.0 81.4 89.7
   –– Fixed-term Contract 0.4 0.8 0.8 100.0 100.0 100.0
Workplace Attributes
   –– Social Environment 39.0 44.0 40.8 80.8 74.4 80.4
   –– Good Colleagues 20.4 26.0 28.2 77.2 60.6 55.8
   –– Possibility for Remote Work 0.3 0.5 0.5 100.0 100.0 100.0
   –– Shared Office Space 0.1 0.0 0.0 100.0
   –– Inclusive Work-life Scheme 6.7 7.2 6.0 98.2 79.3 95.8
Task-Related Attributes
   –– Interesting Tasks 32.4 38.2 22.2 82.5 56.2 96.6
   –– Challenging Tasks 26.9 32.0 19.0 85.0 56.2 94.7
   –– Variation in Tasks 10.3 10.2 15.8 87.5 65.9 42.9
   –– Responsibility in Job 1.5 1.8 10.2 90.5 71.4 12.2
   –– Independence in Performing Tasks 2.4 2.8 5.0 95.8 63.6 35.0
   –– Involves Leadership Responsibility 11.1 13.5 9.0 87.0 35.2 52.8
   –– Work Involves Travelling 7.8 10.2 3.5 93.2 34.1 100.0
Minor Perks
   –– Beautiful Location 0.7 2.2 4.8 96.5 77.8 36.8
   –– Central Location 33.0 37.5 5.2 65.8 11.3 81.0
   –– Company Gym or Sports Team 3.5 5.0 6.5 95.5 70.0 53.8
   –– Parking Space On Premises 0.5 0.5 0.2 99.8 50.0 100.0
   –– Company Vehicle 1.2 1.5 1.0 99.5 66.7 100.0
   –– Any Welfare Scheme 4.0 4.8 2.5 97.2 47.4 90.0
   –– Company Cabin 1.7 2.0 3.2 98.8 100.0 61.5
   –– Occupational Health Service 1.3 1.2 1.2 100.0 100.0 100.0
   –– Company Canteen 0.3 0.8 0.8 99.0 33.3 33.3
   –– Flexible/Extended Holidays 0.4 0.0 0.0 100.0
Number of Job Ads 915,790 400 400 400
Note: See details in the notes to Table 3.
Table A.5: Association Between Posted Employer Values and Vacancy Duration.
Outcome: Log Vacancy Duration
By Gender: By Education:
Pooled Sample Men Women Non-College College
(1) (2) (3) (4) (5)
Panel A: Overall Employer Value
Posted Overall Value ($\tilde{V}_{j}$) -0.154∗∗∗ -0.108∗∗∗ -0.136∗∗∗ -0.121∗∗∗ -0.122∗∗∗
(0.027) (0.025) (0.023) (0.017) (0.018)
Panel B: Pay, Non-Pay, and Job Security Values
Posted Pay Premium ($\tilde{\psi}_{j}$) -0.402∗∗ -0.761∗∗∗ -0.363 -0.378 -0.549∗∗
(0.184) (0.211) (0.187) (0.207) (0.268)
Posted Non-Pay Value ($\tilde{a}_{j}$) -0.573∗∗∗ -0.552∗∗∗ -0.722∗∗∗ -0.430∗∗∗ -0.810∗∗∗
(0.159) (0.173) (0.132) (0.145) (0.175)
Posted Job Security Value ($\tilde{s}_{j}$) -0.086∗∗∗ -0.096∗∗∗ -$0∗∗$082∗∗∗ -0.078∗∗∗
(0.018) (0.018) (0.015) (0.015) (0.021)
Controls:
Occupation Fixed Effects
Industry Fixed Effects
Location Fixed Effects
Observations:
Employer Clusters 1,559 858 850 937 761
Establishments 78,104 42,921 42,479 46,834 38,103
Ads 624,875 468,306 507,506 478,109 499,423
Worker-Years 10,974,238 4,931,815 4,956,647 4,973,189 4,664,927
Robust standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Note: This table documents associations between posted employer values and vacancy duration. Vacancy duration is the number of days between the dates when the ad is posted and unlisted, censored between 1 and 90 days to avoid outliers, averaged at the cluster level before taking logs. We compute the “posted” employer value $(\tilde{V}_{j},\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) associated with each ad as the prediction from regressions of employer values ($V_{j},\psi_{j},a_{j}$, $s_{j}$) on indicators for posted ad attributes. The “posted” employer pay premium $(\tilde{\psi}_{j}$) and non-pay (amenity) value $(\tilde{a}_{j}$) are measured in logs, so the estimated coefficient for each of these can be interpreted as an elasticity, while the overall employer value $(\tilde{V}_{j}$) and job security ($\tilde{s}_{j}$) are scaled by their standard deviation, so the estimated coefficient for each of these represents a standard deviation increase in employer value. Column (1) shows estimates for the “posted” employer values estimated using the pooled sample, while columns (2)-(5) show the corresponding estimates for each population group. The specification in Panel A focuses only on the overall employer value $(\tilde{V}_{j}$), while Panel B includes its components $(\tilde{\psi}_{j},\tilde{a}_{j}$, $\tilde{s}_{j}$) jointly in the same estimation. Estimations are done at the employer-cluster level for the five-period from 2015 to 2019, and are weighted by the number of worker-years in each cluster and population group. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm (Oslo as an own group).
Figure A.1: Sample Vacancy Text: Teacher Substitute.
Substitute in a professional position: One year of temporary work in a 100% position from 01.08.2016 - 31.07.2017. The position entails travel activities in the region and the applicant must have their own car. We are looking for an employee with: - Special needs teacher with competence at the Master’s or Bachelor’s level. - Professional experience from kindergarten/school and relevant skills. - Good collaboration skills, flexibility, and personal suitability. The following applies to the position: - Challenging tasks. - Good working environment in an interdisciplinary team. - Appointment and working conditions according to current agreements and regulations. - Wages in accordance with the agreement. - Compulsory membership in KLP (public sector occupational pension scheme). […]
Note: Translation from Norwegian by the authors.
Figure A.2: Sample Vacancy Text: Civil Engineer.
[…] We need to increase capacity within water- and environmental engineering, and are therefore looking for a civil engineer/engineer with experience within water supply, sewage and urban drainage systems. The position involves exciting and challenging tasks in an interdisciplinary environment. The work will include planning, engineering, and follow-up of private and municipal urban drainage-, water-, and sewage systems. Desired competence: - Engineer / civil engineer - Relevant experience - Good ability to work with colleagues, partners, and clients - Good knowledge of written and spoken Norwegian - New graduates with good results may also be relevant Preferred abilities: - Ability to work independently as well as in a team - Solution-oriented and accurate - Resourceful - Positive and persistent We offer: - An independent position in a solid company in growth and development - Professionally inspiring with talented colleagues - A pleasant and orderly workplace with bright and open premises in central Oslo - Competitive conditions […]
Note: Translation from Norwegian by the authors.
Figure A.3: The Prevalence of Job Attributes Advertised in Vacancy Texts: 2002-2019.
Note: This figure documents the prevalence of job attributes detected in job ads posted in Norway between 2002 and 2019 [N=2,813,645]. The light blue bars show the share of ads detected having each of the distinct job attributes. The dark blue bars show the share of ads detected with at least one attribute within ten broad categories.
Figure A.4: Explained Variation in Publicly Advertised Job Attributes: 2002-2019.
Note: This figure shows pseudo $R^{2}$ [McFadden, 1974] from separate logistic regressions of binary job attributes detected in job ads posted in Norway between 2002 and 2019 [N=2,813,645], on fixed effects denoting unique combinations of 2-digit industries (90 groups), location indicators (10 groups), 2-digit occupations (43 groups), and 200,220 establishments. We start by including industry fixed effects, continue by including industry×$\times$×location fixed effects, and so on. The last set of regressions controls for 563,170 unique combinations. Location indicators group 422 municipalities into ten groups based on the number of workers such that municipalities assigned to the same group have a similar number of workers, with a specific indicator for job postings in Oslo.
Figure A.5: Publicly Advertised Job Attributes By Employer Overall Values.
Note: This figure shows parameter estimates from separate regressions of the share of each ad attribute on overall employer values ($V_{j}$), capturing the change in the fraction of ads with each job attribute associated with a one standard deviation increase in employer value. Estimation are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster. We control for occupation and industry using two-digit occupation and two-digit industry fixed effects and location based on deciles of the number of workers in the municipality of the firm.
Figure A.6: Decomposing the Predictive Power of Publicly Advertised Job Attributes.
(a) Overall Value ($V_{j}$)
(b) Pay ($\psi_{j}$)
(c) Non-Pay ($a_{j}$)
Note: This figure documents adjusted $R^{2}$s from regressions of estimated overall ($V_{j}$), pay ($\psi_{j}$) and non-pay ($a_{j}$) employer values on pay and non-pay job attributes and other ad characteristics detected in the text analysis as in Figure 6, decomposing the overall $R^{2}$s into ten broad categories of pay and non-pay attributes. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, and are weighted by the number of worker-years in each cluster.
Figure A.7: Predictive Power of Publicly Advertised Job Attributes Over Time.
(a) 2000-2004
(b) 2005-2009
(c) 2010-2014
(d) 2015-2019
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 7. Estimations are done at the employer-cluster level for each of the four five-year period from 2000 to 2019, and are weighted by the total number of worker-years in each cluster and period. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
Figure A.8: Heterogeneity in the Predictive Power of Publicly Advertised Job Attributes.
(a) Men
(b) Women
(c) College
(d) Non-college
Note: This figure documents adjusted $R^{2}$ from regressions of estimated population group-specific model parameters on ad attributes and characteristics in the text analysis as in Figure 7, with fixed weights across sample splits. Estimations are done at the employer-cluster level for the five-year period from 2015 to 2019, separately by gender (Panels A-B) and education (Panels C-D), and are weighted by the total number of worker-years in each cluster, as opposed to by cluster and sub-sample as in Figure 7. Overall $R^{2}$s are from regressions that include ad attributes and other ad characteristics, while partial $R^{2}$s are from regressions that also control for the composition of industry, occupation and location in each cluster. Industry and occupation controls are defined at the two digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having similar number of workers (Oslo as an own group).
Appendix B Additional Details on the Text Analysis
Section B.1 documents how we clean the text of job ads before detecting job attributes. Section B.2 describes one of the processes we use to generate the complete lists of attributes we search for in vacancy texts. Section B.3 provides descriptions of detailed attributes and displays some of the phrases used to detect these attributes.
B.1 Text Cleaning of Job Ads
Before attempting to extract attributes from job ads we start with some simple cleaning of the texts contained in vacancies. We start by searching for HTML tags in the vacancies. We do this to avoid considering words that are part of HTML code to indicate job attributes, and to reduce the number of words considered by the model we use to generate phrases for attribute detection. HTML tags always start with “$<$” and end with “$>$”. We keep a record of which ads are detected with this pattern and remove the corresponding tags.
Second, we replace numbers in the texts with flags indicating their range and whether the number was followed by a percentage. We do this to be able to group similar phrases containing different numbers. For example, “50 %” and “40 %” are commonly used to indicate part time jobs. By grouping every phrase of the form ‘{a number below 100} %’ we avoid having to search for every number less than a hundred followed by a percentage sign. We start by flagging numbers less than 100 (equal to 100) followed by a percentage sign or the word “percentage”. We then create separate flags for each remaining integer less than or equal to four, remaining numbers between 0 and 100, the number 100, and numbers greater than 100.
Before storing the vacancies, we split the ads into distinct sentences, and sentences into distinct words. This is primarily done because the language model we use to expand the dictionary of target phrases expects input in this form. Splitting sentences also has the advantage of separating the words at the end of one sentence from the words in the beginning of the next. We use the Python package nltk to do these splits. The package uses a model that is trained on a corpus in Norwegian, and it avoids splitting at periods that are not at the end of sentences, e.g., as part of a common abbreviation.
B.2 Extracting Common Attributes from “We Offer” Lists
Our choice of which pay and non-pay attributes to include in the text analysis is based partly on commonly advertised attributes extracted from “we offer” sections of job ads. This section documents this process, and shows that attributes related to social environment, on-the-job training, tasks, pension schemes, and insurance schemes are commonly advertised in “we offer” sections of publicly posted vacancies. Our approach has four broad steps: (i) we extract lists from the corpus of job ads; (ii) we sort the lists into four distinct topics using a common model from natural language processing; (iii) we identify the topic that contains “we offer” sections by inspecting words and lists that, according to the model, are closely associated with the “we offer” sections; and (iv) we extract common phrases from the lists that are sorted into the “we offer” category. We assign some of these phrases to attribute categories, and add these categories to the list of attributes to search for in vacancies.
Extracting lists from vacancies
To identify “we offer” sections in vacancies, we start by identifying and extracting a large collection of lists from the full corpus of job ads. Similar information in job ads is commonly grouped in lists, such as skill requirements, tasks, and the pay and non-pay content (i.e., “we offer” section). We extract lists by flagging consecutive instances of the html tag “<li>”, which is used to indicate list elements in web pages, and plain text lines that start with a centered dot (·), hyphen (-) or a star (*). We combine consecutive sentences with this flag with the previous sentence (the list header) into separate observations. This process allows us to extract sections of texts that are likely to center on a single topic like skill requirements, tasks and pay and non-pay attributes. Although it possibly omits some lists, we are able to extract close to 1.5 million lists from around 50 percent of the vacancies.
Sorting lists into different topics
To distinguish “we offer” lists from lists covering other topics, we apply Latent Dirichlet Allocation (LDA) to the collection of lists [Blei et al., 2003]. Latent Dirichlet Allocation is commonly used in natural language processing to partition a collection of texts into different topics. LDA partitions the texts into $K$ categories such that words that commonly occur in topic $k$ are unlikely to occur in other topics. The model performs well when each topic is associated with words that are not frequently used in other topics. However, the model ignores how words are ordered, making the individual words that are used to assign texts to topics less informative. For example, the model splits the phrase “good colleagues” into two observations, “good” and “colleagues.” While “good” is commonly used to indicate workplace amenities or pay, it does not contain information about the specific attribute without its context. Additionally, some lists combine “we offer” sections with other topics, such as skill requirements. We therefore expect to lose some of the “we offer” content that is assigned to other topics by the model.
Latent Dirichlet Allocation partitions texts into topics by modelling each individual word as being drawn from a mixture model: Each word is drawn from one of $K$ word-distributions, each corresponding to a different topic. The texts (lists) are associated with separate mixtures over the $K$ topics, meaning that the probability that a word is drawn from topic $k$ varies across texts (lists). That is the probability that the word $i$ in text $j$ ($w_{ij}$) is the word “offer” is
$\Pr(w_{ij}=\textnormal{``offer''})=\sum_{k=1}^{K}\Pr(w_{ij}=\textnormal{``offer''}|\textnormal{topic}_{ij}=k)\Pr(\textnormal{topic}_{ij}=k)$
We set the number of topics to 4 and estimate the model with maximum likelihood using the implementation provided in the Python package gensim.
Labelling topics
After estimating the model, we label the topics into four identified topics by considering the word distributions and the text (list) specific topic mixtures. Table B.1 documents the ten words with the highest association with each topic, i.e., the highest probability of being drawn in each topic. Words in topics 1-3 are associated with 1) (both pay and non-pay) job amenities, 2) skill requirements, and 3) task descriptions, while words in topic 4 contains a mixture of requirements (e.g. education and police certificate), additional information (e.g. contact information), and ambiguous words (numbers and “work”). We label the topics 1-3 according to these observations, and topic 4 as a “Residual Category”.
The “we offer” topic contains words commonly used to indicate a good social environment, information about pay, and words used to describe task related attributes that might be attractive to workers (e.g., challenging tasks). We note that some of these words are ambiguous outside of their original context, such as “development”, which can be used to indicate personal development or tasks that includes development of projects.
We validate these labels by inspecting the lists with the highest association with each topic in Figure B.1-B.4. These lists serve as stylized examples of each topic and omit only considering one word at the time. These lists largely confirms the labels associated with each topic. The list most strongly associated with the “We Offe”-topic starts with the header “we offer:” and contains information about various non-pay amenities. The list most strongly associated with the ”Residual Category”-topic contains various information about the job, such as the hiring process and and information about pay. Overall, these lists suggests that lists in the “We Offer” category contain dense information about amenities, although we are likely to lose some information about pay and non-pay in lists sorted into other categories (e.g., information about pay in lists assigned to the “Residual Category”).
Extracting common phrases from “We Offer” sections
To identify specific job attributes commonly advertised in vacancies, we now focus on common words and phrases in lists associated with the “we offer” topic identified by the model. We label each list with the topic that has the most mass in the list’s associated topic distribution, and extract about 260,000 lists within the “we offer” label (i.e., 17 percent of all lists).
Next, we extract and inspect the 200 most common unigrams, bigrams and trigrams, excluding phrases that includes stopwords (e.g., “and”). We add phrases that indicate (pay or non-pay) job amenities to our dictionary of phrases, and create new attribute categories whenever a phrase indicates an attribute not currently in our list.
Table B.2 documents the 10 most common unigrams, bigrams, and trigrams in “we offer” sections, excluding phrases without a clear reference to a particular job attribute. The table documents that these sections commonly refer to the social environment and pay-related attributes (pay, pension and insurance scheme, and on-the-job training). This procedure does not uncover many attributes related to hours of work, which we suspect is referred to in other sections of ads (e.g., practical information in the “Residual Category”).
Table B.1: Words Used to Separate Lists Into Topics
Topic 1: We-Offer Topic 2: Skill Requirements Topic 3: Tasks Topic 4: Residual Category
Word Probability Word Probability Word Probability Word Probability
(1) (2) (3) (4) (5) (6) (7) (8)
“offer” 0.037 “experience” 0.038 “tasks” 0.025 Number $>99$ 0.044
“good (plural)” 0.032 “good (plural)” 0.027 “responsibility” 0.011 Number in [5-99] 0.024
“workplace environment” 0.029 “good” 0.024 “to follow up” 0.010 “have to” 0.016
“good (alternative form)” 0.026 “qualifications” 0.021 “contribute” 0.009 “Field of education” 0.009
“pay” 0.019 “ability” 0.016 “cooperation” 0.009 “police certificate” 0.008
“professional” 0.017 “Norwegian” 0.014 “other” 0.008 “contact information” 0.007
“exciting” 0.015 “relevant” 0.014 “tasks” 0.007 “level of education” 0.007
“tasks” 0.014 “written” 0.014 “participate” 0.007 “The position” 0.007
“developement” 0.014 “oral” 0.013 “customers” 0.006 “work” 0.006
“challenging” 0.010 “traits” 0.013 “developement” 0.005 “two” 0.006
Note: This table documents the words that are most strongly associated with the topics identified by applying Latent Dirichlet Allocation to our collection of lists. Columns (2), (4), (6), and (8) contains estimated probabilities of drawing the corresponding word given that the word is drawn from the distribution associated with that topic. The table shows the words with the 10 largest probabilities for each topic. “Number $>99$” and “Number in [5-99]” are flags that we have inserted into the vacancy in the cleaning step to group numbers by range. Translation from Norwegian by the authors. Each entry is a single Norwegian word, although the English translations sometimes require more than one word. Clarifications in parentheses are added whenever multiple distinct Norwegian words have identical English translations.
Figure B.1: Sample List Assigned to Topic “We Offer”.
We offer: - a unique opportunity to set foot in Norway’s leading exercise company. - an incredibly exciting job that offers some real challenges that provide a lot of development as a person. - very competitive conditions. - we offer the industry’s best training program and very good career opportunities for the right person.
Note: This figure shows the list that the model most strongly associates with the “we offer” topic. This means that the models assigns a probability of almost 1 for words in this list to be drawn from the distribution associated with “we offer.” Translation from Norwegian by the authors.
Figure B.2: Sample List Assigned to Topic “Skill Requirements”.
We are looking for someone who: - has relevant education in economics, marketing and/or technical subjects. - has good knowledge of the Norwegian retail. - has ability to motivate and inspire others. - communicates effectively and well, written and oral. […]
Note: This figure shows the list that the model most strongly associates with the “skill requirement” topic. This means that the models assigns a probability of almost 1 for words in this list to be drawn from the distribution associated with “skill requirement.” Translation from Norwegian by the authors.
Figure B.3: Sample List Assigned to Topic “Tasks”.
Tasks: - personnel responsibility for all resources associated with the section. - financial management and responsibility for results within the area. - coordination with other sections and departments. - be the driving force behind the company’s proactive approach to security and risk. […]
Note: This figure shows the list that the model most strongly associates with the topic “tasks”. This means that the models assigns a probability of almost 1 for words in this list to be drawn from the distribution associated with the topic “tasks.” Translation from Norwegian by the authors.
Figure B.4: Sample List Assigned to Topic “Residual Category”.
Start date, salary placement, and general information about the position: - the position is a temporary position with a duration until March and is available for immediate entry. - the position of chief police officer is placed in salary plan [number $>100$] to [number $>100$]. - our personnel policy is to have a employees from different backgrounds. We therefore encourage women and people with minority background to apply. - the statutory pension premium is deducted from the salary - spotless conduct is required. A certificate from the police will be obtained. - applicants may be called to an interview. […]
Note: This figure shows the list that the model most strongly associates with the residual category. This means that the models assigns a probability of almost 1 for words in this list to be drawn from the distribution associated with the residual category. Translation from Norwegian by the authors.
Table B.2: Most Frequent Phrases in “We-Offer” Lists.
Rank a Phrases a English Translation Occurrences a Assigned Attribute Category
(1) (2) (3) (4) (5)
Panel A: Unigrams
3 “arbeidsmiljø” “workplace environment” 138,459 Social Environment
5 “lønn” “pay” 91,477 Any Other Mention of Pay
6 “spennende” “exciting” 88,695 Interesting Tasks
12 “utfordrende” “challenging” 59,382 Challenging Tasks
21 “opplæring” “training” 37,351 On-the-Job Training
24 “pensjons” “pension” 32,520 Pension Scheme
25 “pensjonsordning” “pension scheme” 32,434 Pension Scheme
26 “utviklingsmuligheter” “opportunities for development” 31,724 Good Career Paths
28 “forsikringsordninger” “insurance schemes” 30,366 Insurance Scheme
36 “oslo” “oslo” 23,336 Central Location
Panel B: Bigrams
1 “godt arbeidsmiljø” “good workplace environment” 65,143 Social Environment
2 “konkurransedyktige betingelser” “competitive conditions” 38,016 Competitive Pay
3 “gode pensjons” “good pension” 22,344 Pension Scheme
4 “faglig utvikling” “professional development” 18,415 On-the-Job Traning
5 “varierte arbeidsoppgaver” “varying tasks” 18,085 Variation in Tasks
7 “utfordrende arbeidsoppgaver” “challenging tasks” 17,777 Challenging Tasks
8 “fleksibel arbeidstid” “flexible work hours” 16,700 Possibility to Work Flexible Hours
10 “personlig utvikling” “personal development” 15,617 On-the-Job Traning
11 “trivelig arbeidsmiljø” “pleasant working environment” 14,591 Social Environment
12 “god pensjonsordning” “good pension scheme” 12,854 Pension Scheme
Panel C: Trigrams
1 “høyt faglig nivå” “high professional level” 6,021 Good Colleagues
2 “god pensjonsordning gjennom” “good pension scheme through ” 3,716 Pension Scheme
7 “godt faglig miljø” “good professional environment” 2,134 Good Colleagues
8 “pensjonsordning gjennom klp” “pension scheme through klp” 2,009 Pension Scheme
9 “godt sosialt miljø” “good social environment” 1,905 Social Environment
10 “gjennom statens pensjonskasse” “public pension fund” 1,805 Pension Scheme
17 “faglig dyktige kollegaer” “professionaly skilles colleagues ” 1,195 Good Colleagues
18 “pensjonsordning gjennom statens” “pension scheme through public” 1,183 Pension Scheme
19 “ulykkesforsikring samt fritidsulykkeforsikring” “sport and general accident insurance” 1,159 Insurance Scheme
20 “svært godt arbeidsmiljø” “very good social environment” 1,129 Social Environment
Note: This table documents the words occurring most frequently in the “we offer” lists extracted from the vacancies. Column 3 shows which attribute category we have assigned the word to. These categories are added to the list of attributes we search for in the vacancies. Not every word is assigned to an attribute category. The translations are provided by the authors. Every phrase is originally one word in Norwegian.
B.3 Description of Detailed Pay and Non-Pay attributes
Tables B.3 to B.5 provides descriptions of each of the attribute we search for in job ads, as well as examples of phrases we use to detect these attributes.
Table B.3: Description of Pay Related Attributes.
Attribute Description Example Phrases (translated) Example Phrases (Norwegian)
Compensation Scheme:
– Compensation Level Indication of pay, pay level, or pay range (e.g., steps in collective bargaining agreements). wage step, wage range, wage rate lønnstrinn, lønnsspenn, lønnssats
– Competitive Pay Indication of good or competative salary. good pay conditions, competitive wage, favorable pay conditions gode lønnsbetingelser, konkurransedyktig lønn, gunstige lønnsbetingelser
– Collective Agreement Pay Explicit reference to a collective bargaining agreement between trade unions and employers’ associations. wage step, tariff agreement, the main tariff agreement [public sector] lønnstrinn, tariffavtale, hovedtariffavtalen
– Incentive Pay Scheme Pay scheme dependents on an individual worker’s or company’s performance. commission, bonus scheme, sales bonus provisjon, bonusordning, salgsbonus
– Hiring Bonus One time bonus at time of hiring, usually to encourage applicants. recruitment supplement, establishment supplement rekrutteringstillegg, rekrutteringstilskudd
– Good Overtime Pay Reference to the level of overtime pay. overtime pay, paid overtime overtidsbetalt, betalt overtid
– Any Other Mention of Pay Any other reference to pay, excluding any of the attributes listed above. pay, income, fixed salary lønn, inntekt, fastlønn
Financial Attributes:
– Pension Scheme Position covered by a pension scheme. Includes reference to mandatory and additional pension schemes. pension scheme, occupational pension scheme, contribusion based pension savings pensionsordning, tjenestepensjonsordning, inskuddspensjon
– Insurance Scheme Employer-sponsored insurance. Includes mandatory and additional insurance schemes, and coverage of proparty damage, various health insurances and similar. insurance scheme, group life insurance, life insurance scheme, disability insurance, health insurance forsikriningsordning, gruppelivsforsikring, livsforsikring, uføreforsikring, helseforsikring
– Mortgage Possibility Employer provides access to mortgage schemes with favorable conditions. mortgage scheme, favorable mortgage boliglånsordning, gunstig lånetilbud
Career Opportunities:
– Good Career Paths Suggests that the position is a good step in advancing a career, or that there are good opportunities to advance within the employer. career opportunities, promotions, internal career karrieremuligheter, forfremmelse, intern karriere
– On-the-Job Training The position includes a training program or internship or opportunities to learn new skills. training program, professional developement, internship etterutdaningsmuligheter, faglig utvikling, praksisplass
Note: This table describes each of the pay-related attributes detected in vacancies and documents some of the phrases used to detect these attributes. The phrases in Norwegian are used in the analysis, and the translations to English are for illustration. Comments to individual phrases are displayed in brackets.
Table B.4: Description of Non-Pay Related Attributes (Part 1).
Attribute Description Example Phrases (translated) Example Phrases (Norwegian)
Hours of Work:
– Full-time Contract Full time position, usually 37.5 hours per week. full-time position, full-time employee, 100% fulltidsstilling, fulltidsansatt, 100%
– Part-time Contract Part-time position. part time, part-time position, 50% [includes all numbers less than 100] deltidsstilling, deltidsjobb, 50%
– Full-time/Part-time Choice Employer open for both full-time and part-time hires. full-time or part-time, full-time/part-time heltid eller deltid, fulltid/deltid
– On-call Employment Positions where the employer needs additional on-call employees, often hired temporarily. on-call help, on-call substitute, calling help [norwegian expression], ekstrahjelp, tilkallingsvikar, ringevikar
Convenient Hours:
– Possibility to Work Flexible Hours Flexible work hours, commonly reflecting workers’ discretion to choose when to start or stop work during a shift. flexible working hours, flexible hours scheme fleksibel arbeidstid, fleksibel arbeidstidsordning
– Regular Daytime Work Schedule Regular, day-time work hours, i.e. working on evenings, nights or weekends is not required. regular working hours normal arbeidsdag
– Exempt from Work Hour Regulations The job is exempt from the work hours regulation covering most workers, usually leaders, academic staff or similar. exempt from the working environment act, leader position [legal term], doctoral fellow [always exempted] untatt arbeidsmiljøloven, ledende stilling, doktorgradsstilling
Inconvenient Hours:
– Shift Work Position requires shift work where employees usually switch between working day, evening, and night shifts. shift work, rotation scheme, night shift skiftarbeid, turnusordning, natturnus
– Weekend/Evening/Night Work Position requires shift work involving either working on weekends, evening, and/or night. night shift, evening work,late night shift nattursnus, kveldsarbeid, senvakt
– Overtime Work Required Indication that some overtime work is required in the position. overtime work, some overtime expected, overtime in periods overtidsarbeid, påregnes noe overtid, overtid i perioder
Contract Duration:
– Permanent Job Permanent job. permanent position, permanent employment fast stilling, fast ansettelse
– Temporary Job Temporary job, including substitute positions. temporary job, substitute position, one year substitute midlertidig stilling, vikariat, ettårsvikariat
– Fixed-term Contract This position is restricted to a fixed period, and is exempted from Norwegian restriction of the length of temporary spells. fixed-term contract [legel term] åremål
Workplace Attributes:
– Social Environment Indication of good social environment. good social environment, good working community, great colleagues godt arbeidsmiljø, godt arbeidsfellessakap, flotte kolleger
– Good Colleagues Professional, knowledgable, helpful, or in other ways good colleagues. skilled colleagues, good colleagues, knowledgeable colleagues, helpful colleagues dyktige kolleger, gode kolleger, kunnskapsrike kolleger, hjelpsomme kolleger
– Possibility for Remote Work Possibility to work remotely. home office, home office scheme, work from home hjemmekontor, hjemmekontorløsning, arbeide hjemmefra
– Shared Office Space Shared office space. shared office space, open office space kontorlandskap, åpent landskap
– Inclusive Work-life Scheme Indication that the employer provides an inclusive work-life scheme, which is a formal work-life scheme that provides more generous sick-leave arrangements. inclusive work-life scheme, IA [abbrevation] inkluderende arbeidsliv, IA
Note: This table and the next describe each of the non-pay-related attributes detected in vacancies and document some of the phrases used to detect these attributes. The phrases in Norwegian are used in the analysis, and the translations to English are for illustration. Comments to individual phrases are displayed in brackets.
Table B.5: Description of Non-Pay Related Attributes (Part 2).
Attribute Description Example Phrases (translated) Example Phrases (Norwegian)
Task-Related Attributes:
– Interesting Tasks Indicates interesting, exciting, or meaningful tasks. Includes description of company, sector, or industry. interesting, exciting, meaningful interessante, spennende, meninsfylt
– Challenging Tasks Indicates challenging tasks. challenging, demanding utfordrende, krevende
– Variation in Tasks Varying, nonmonotonous tasks. varied tasks, versatile tasks varierte arbeidsoppgaver, allsidige oppgaver
– Responsibility in Job Involves responsibilities. Includes responsibility as leader or project manager. full responsibility, a lot of responsibilities, leadership responsibilities totalansvar, mye ansvar, lederansvar
– Independence in Performing Tasks Freedom to choose how to approach and solve tasks. manage the working day, influence your own, shape your own styre arbeidsdagen, påvirke egen, utforme egen
– Involves Leadership Responsibility The job involves leadership responsibility. looking for CEO, leadership role, communications director søker daglig leder, lederrolle, kommunikasjonsdirektør
– Work Involves Travelling The job requires traveling. business trips, travel days, travel activity arbeidsreiser, reisedøgn, reiseaktivitet
Minor Perks:
– Beautiful Location Describes the job location’s environment as beautiful. beautiful nature, magnificent nature, great tracking environment flott natur, praktfull natur, flott turterreng
– Central Location Central location, including mentions of major cities. oslo [capial], centrally located, close to the city center oslo, sentral beliggenhet, sentrumsnært
– Company Gym or Sports Team The company offers the opportunity for paid exercise during work hours, access to fitness center/equipment, and/or has a sports team or the like. company sports team, exercise during work-hours, fitness room, fitness center membership bedriftsidrettslag, trening i arbeidstiden, treningsrom, treningsavtale
– Parking Space On Premises Available/free parking spaces on premises. parking facilities, free parking parkeringsmuligheter, gratis parkering
– Company Vehicle Access to company vehicle during work hours, sometimes involves company vehicle available full-time. company vehicle, company vechicle scheme, leasing vechicle firmabil, bilordning, leasingbil
– Any Welfare Scheme Generic description of company welfare scheme. welfare schemes, personnel schemes, employee benefits velferdsordninger, personalordninger, personalgoder
– Company Cabin The company has access to a cabin usable for employees. These cabins can be accessable for company arrangements or for employees’ access during holidays or vacations. company cabins, personnell cabin, vacation apartments firmahytte, personalhytte, ferieleiligheter
– Occupational Health Service Access to health professionals and doctors. Involves preventions and treatment of injuries/sickness. occupational health service, company doctor bedriftshelsetjeneste, bedriftslege
– Company Canteen Access to canteen. canteen scheme, personnell canteen kantineordning, personalkantine
– Flexible/Extended Holidays Description of the total length of vacation, or of extended length compared to mandatory vacation. x week vacation [where x is a number], extra vacation x ukers ferie, ekstra feie
Note: This table and the previous describe each of the non-pay-related attributes detected in vacancies and document some of the phrases used to detect these attributes. The phrases in Norwegian are used in the analysis, and the translations to English are for illustration. Comments to individual phrases are displayed in brackets.
Appendix C Additional Details on the Model
Here we describe the transitions to and from non-employment in our structural model. The flow measure of workers from employer $j$ to non-employment $N$ is the sum of a relocation part $M_{jN}^{R}$ and a voluntary part $M_{jN}^{V}$. The relocation flow to non-employment is simply the fraction of employment at firm $j$ hit by an exogenous separation shock $\delta_{j}$
$M_{jN}^{R}=L_{j}\cdot\delta_{j}.$ (B.1)
The voluntary flow to non-employment is given by the fraction of workers who decide to leave employer $j$
$M_{jN}^{V}=L_{j}\cdot(1-\delta_{j}-\rho_{j})\cdot(1-\lambda_{1})\cdot\text{Pr}(N\succ j),$ (B.2)
where $1-\delta_{j}-\rho_{j}$ is the fraction of workers who are not forced to reallocate and $1-\lambda_{1}$ is the fraction of workers who do not have an alternative offer in the current period. Similarly to the case of voluntary flows between employers (2), the probability that a worker employed at $j$ moves to non-employment is
$\text{Pr}(N\succ j)=\text{Pr}(V_{N}+\varepsilon_{N}\geq V_{j}+\varepsilon_{j})=\frac{\exp(V_{N})}{\exp(V_{j})+\exp(V_{N})},$ (B.3)
given the maintained assumption of i.i.d. idiosyncratic taste shocks $(\varepsilon_{j},\varepsilon_{N})$ from a Extreme Value Type I distribution.
All transitions from non-employment $N$ to employer $j$ are voluntary. The decision to accept a job opportunity is again subject to i.i.d. taste shocks $(\varepsilon_{N},\varepsilon_{j})$. The flow measure of workers moving from non-employment to employer $j$ is given by
$M_{Nj}=L_{N}\cdot\lambda_{0}\cdot f_{j}\cdot\frac{\exp(V_{j})}{\exp(V_{N})+\exp(V_{j})}.$ (B.4)
In Equation (B.4), $L_{N}$ is the measure of workers in non-employment, $\lambda_{0}$ is the probability that a non-employed worker gets an offer, and $f_{j}$ is the chance the probability that offer is from employer $j$. For the identification result, it is required that the offer distribution $f$ is the same for employed and non-employed worker, which we assume.
With the notation introduced in Equation (B.4), the value function of non-employment follows directly as
$V_{N}=u_{N}+\beta\left\{\lambda_{0}\sum_{k}f_{k}\mathbf{E}\big{[}\max\big{\{}V_{k}+\varepsilon_{k},V_{N}+\varepsilon_{N}\big{\}}\big{]}+(1-\lambda_{0})\mathbf{E}\big{[}V_{N}+\varepsilon_{N}\big{]}\right\},$ (B.5)
where $u_{N}$ is the flow value of non-employment and $\beta$ the discount factor. The continuation value is made of two terms. With probability $\lambda_{0}$, non-employed workers get a draw from the offer distribution $f$, which they decide to accept or reject given their draw of taste shocks. With probability $\lambda_{0}$, they remain in non-employment. All expectations $\mathbf{E}[.]$ are taken over the i.i.d. taste shocks $(\varepsilon_{k},\varepsilon_{N})$.
Appendix D Predictive Power of Publicly Posted Job Attributes
Which job attributes contain the most predictive power? In Section 5.2 we decompose the overall $R^{2}$ from regressions of employer values on ad attributes by sequentially adding groups of attributes. This approach abstracts away from the 47 individual attributes and four other ad characteristics, and depends on the order of which attributes are added. This section documents that the main patterns in Section 5.2 and Appendix Figure A.6 hold when we relax the dependency on the ordering and focus on individual attributes and ad characteristics.
To do this, we implement a method that sequentially identifies the attributes with the highest predictive power. We start with a regression containing only a constant term, and then add the attribute with the largest adjusted (partial) $R^{2}$ among the 51 attributes as an explanatory variable (with additional controls). We continue by sequentially adding attributes that increase the explanatory power further, adding one variable at a time. This procedure delivers a tractable sequence of explanatory variables where variables with the most explanatory power are added first. It reveals, at each step, the variable with the largest predictive power given the variables already added. However, it does not identify the set of n variables with the largest joint predictive power at each step n.
Figure D.1: Explanatory Power as a Function of the Number of Regressors.
(a) Total Value ($V_{j}$)
(b) Pay ($\psi_{j}$)
(c) Non-Pay ($a_{j}$)
(d) Job Security ($s_{j}$)
Note: This figure documents the change in explanatory power in regressions models when adding variables representing individual job attributes one at a time, in terms of adjusted $R^{2}$ for regressions without additional controls (black line) and in terms of partial $R^{2}$ for regressions with additional controls for industry, occupation and location (blue line). In each case, we start with a regression containing only a constant term, and iteratively add the variable that increases the adjusted $R^{2}$ the most.
Figure D.1 shows that 10-20 attributes contain most of the explanatory power in regressions of each model parameter on job attributes. The black lines show sequential adjusted $R^{2}$ measures in regressions without additional controls, while the blue lines show sequential partial $R^{2}$ measures in regressions with additional controls. Table D.1 summarizes the first 5, 10, and 20 attributes added to these regressions, with and without additional controls.
For regressions without additional controls, we find that attributes related to convenient or inconvenient hours are important predictors (i.e., among the first five added variables) for all model components. This confirms the results in Section 5.2. For pay values, we find that attributes related to financial attributes and hours of work are also important. For non-pay values, we find that “on-the-job training” and “permanent job” are also among the most important predictors, together with “variation in tasks” and the number of words in ads. This confirms the observation that non-pay values are related to different types of attributes posted in job ads (Section 5.2), and uncovers that “shift work” carries a significant portion of the prediction power among inconvenient hours attributes. The pattern for job security is less clear, with no real prediction value from attributes related to contract duration.
For regressions with additional controls, we find a somewhat different pattern. Reassuringly, “permanent job” is the first attribute added in regressions of job security with additional controls. Yet, again, most of the predictive power stems from attributes related to convenient or inconvenient hours, besides financial attributes and hours of work.
Table D.1: The Most Predictive Publicly Advertised Job Attributes.
No Additional Controls (Adjusted $R^{2}$) With Additional Controls (Partial $R^{2}$)
Total Value ($V_{j}$) Pay ($\psi_{j}$) Non-Pay ($a_{j}$) Job Security ($s_{j}$) Total Value ($V_{j}$) Pay ($\psi_{j}$) Non-Pay ($a_{j}$) Job Security ($s_{j}$)
Detected Job Attribute: (1) (2) (3) (4) (5) (6) (7) (8)
Compensation Scheme
– Compensation Level
– Competitive Pay ∗∗ ∗∗
– Collective Agreement Pay ∗∗
– Incentive Pay Scheme ∗∗ ∗∗∗ ∗∗ ∗∗
– Hiring Bonus ∗∗
– Good Overtime Pay
– Any Other Mention of Pay
Financial Attributes
– Pension Scheme ∗∗∗ ∗∗∗
– Insurance Scheme ∗∗∗ ∗∗∗ ∗∗∗
– Mortgage Possibility ∗∗ ∗∗∗
Career Opportunities
– Good Career Paths ∗∗
– On-the-Job Training ∗∗∗ ∗∗∗
Hours of Work
– Full-time Contract ∗∗
– Part-time Contract ∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Full-time/Part-time Choice ∗∗∗ ∗∗ ∗∗ ∗∗∗ ∗∗
– On-call Employment
Convenient Hours
– Possibility to Work Flexible Hours ∗∗ ∗∗ ∗∗∗ ∗∗∗
– Regular Daytime Work Schedule ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗
– Exempt from Work Hour Regulations ∗∗∗ ∗∗∗
Inconvenient Hours
– Shift Work ∗∗∗ ∗∗∗ ∗∗∗ ∗∗∗ ∗∗ ∗∗
– Weekend/Evening/Night Work ∗∗∗ ∗∗ ∗∗
– Overtime Work Required
Contract Duration
– Permanent Job ∗∗ ∗∗∗ ∗∗ ∗∗∗
– Temporary Job
– Fixed-term Contract
Workplace Attributes
– Social Environment ∗∗
– Good Colleagues ∗∗∗
– Possibility for Remote Work ∗∗ ∗∗
– Shared Office Space
– Inclusive Work-life Scheme
Task-Related Attributes
– Interesting Tasks ∗∗
– Challenging Tasks
– Variation in Tasks ∗∗∗ ∗∗ ∗∗∗
– Responsibility in Job ∗∗
– Independence in Performing Tasks ∗∗ ∗∗
– Involves Leadership Responsibility ∗∗ ∗∗∗
– Work Involves Travelling
Minor Perks
– Beautiful Location
– Central Location ∗∗ ∗∗
– Company Gym or Sports Team ∗∗
– Parking Space On Premises
– Company Vehicle ∗∗
– Any Welfare Scheme ∗∗ ∗∗∗ ∗∗∗
– Company Cabin
– Occupational Health Service
– Company Canteen
– Flexible/Extended Holidays
Other Ad Characteristics
– Number of Words ∗∗ ∗∗∗ ∗∗∗
– Posted Online ∗∗ ∗∗
– Employer Name Disclosed
– Number of Attributes ∗∗
∗∗∗ Model with 5 regressors, ∗∗ Model with 10 regressors, Model with 20 regressors.
Note: This table documents the variables selected in restricted models with 5, 10 and 20 attributes. Columns (1)-(4) show the first variables selected in regressions without including additional controls, while columns (5)-(8) show the first variables selected in regressions including controls for occupation, industry and location. We start with a regression on a constant term, and iteratively add variables that increase the adjusted (partial) $R^{2}$ the most. Regressions with 5 regressors include the first 5 variables added using this procedure. Regression with 10 regressors also includes variables selected in models with 5 regressors, etc. Industry and occupation controls are defined at the two-digit level. Location indicators split local municipalities into deciles, with municipalities in the same group having a similar number of workers.